### v4

# Relative Fractional Independence Number and Its Applications

Abstract
We define the relative fractional independence number of a graph $G$ with respect to another graph $H$, as
$\alpha^{*}(G|H)=\max_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)},$
where the maximum is taken over all graphs $W$, $G\boxtimes W$ is the strong product of $G$ and $W$, and $\alpha$ denotes the independence number. We give a non-trivial linear program to compute $\alpha^{*}(G|H)$ and discuss some of its properties. We show that $\alpha^{*}(G|H)\geq\frac{X(G)}{X(H)}\geq\frac{1}{\alpha^{*}(H|G)},$ where $X(G)$ can be the independence number, the zero-error Shannon capacity, the fractional independence number, the Lovász number, or the Schrijver’s or Szegedy’s variants of the Lovász number of a graph $G$. This inequality is the first explicit non-trivial upper bound on the ratio of the invariants of two arbitrary graphs, as mentioned earlier, which can also be used to obtain upper or lower bounds for these invariants. As explicit applications, we present new upper bounds for the ratio of the zero-error Shannon capacity of two Cayley graphs and compute new lower bounds on the Shannon capacity of certain Johnson graphs (yielding the exact value of their Haemers number). Moreover, we show that $\alpha^{*}(G|H)$ can be used to present a stronger version of the well-known No-Homomorphism Lemma.  [^1] A shortened conference version of this work (containing some of the results) is currently under review for presentation at the 2024 IEEE Information Theory Workshop.
Keywords: Independence number, Zero-error Shannon capacity, Lováz number, Haemers number, Homomorphism
1 Introduction
Let $G$ be a finite, undirected graph without a loop or multiple edges. The set $\mathcal{V}(G)$ denotes the vertex set of $G$, and $\mathcal{E}(G)$ represents the edge set of $G$. An independent set of $G$ is a subset $\mathcal{S}$ of $\mathcal{V}(G)$ such that no two vertices in $\mathcal{S}$ are adjacent in $G$, i.e., $uv\notin\mathcal{E}(G)$ for all $u,v\in\mathcal{S}$. The independence number (or the packing number) of a graph $G$, denoted by $\alpha(G)$, is the size of the biggest independent set of $G$. Computing $\alpha(G)$ is an NP-hard problem [1]. The strong product of two graphs, $G\boxtimes H$, is a graph whose vertex set is the Cartesian product of the vertex sets of $G$ and $H$. Distinct vertices $(u,u^{\prime})$ and $(v,v^{\prime})$ are adjacent in $G\boxtimes H$ iff either $u=v$ and $u^{\prime}v^{\prime}\in\mathcal{E}(H)$, or $uv\in\mathcal{E}(G)$ and $u^{\prime}=v^{\prime}$, or $uv\in\mathcal{E}(G)$ and $u^{\prime}v^{\prime}\in\mathcal{E}(H)$.
The zero-error Shannon capacity (or the Shannon number) of a graph $G$ is defined as
$\mathscr{C}(G)=\lim_{n\rightarrow\infty}\alpha(G^{n})^{\frac{1}{n}},$ (1)
where $G^{n}$ is the strong graph product of $G$ with itself $n$ times. It is known that the above limit exists and furthermore $\mathscr{C}(G)\geq\alpha(G^{n})^{\frac{1}{n}}$ for every $n\geq 1$. This graph invariant was introduced by Shannon in 1956 and gives a measure of the optimal zero-error performance of an associated memory-less communication channel [2].
In general, computing the exact value of $\mathscr{C}(G)$ is a challenging problem, even for simple graphs like a cycle of length $7$, $C_{7}$, $\mathscr{C}(C_{7})$ is not known. Lovász proved that $\mathscr{C}(C_{5})=\sqrt{5}$, where $C_{5}$ is a cycle of length $5$. For more results on the Shannon capacities of odd cycles or their compliments, see [3, 4, 5]. See also [6]. Alon and Lubetzky in [7] showed that the series of independence numbers in strong powers of a fixed graph can exhibit a complex structure, implying that the Shannon capacity of a graph cannot be approximated (up to a sub-polynomial factor of the number of vertices) by any arbitrarily large, yet fixed, prefix of the series of independence numbers in strong powers. Nonetheless, various upper bounds on the Shannon capacity are obtained in the literature. As an example, note that
$\alpha(G^{n})=\prod_{i=1}^{n}\frac{\alpha(G^{i})}{\alpha(G^{i-1})}=\prod_{i=1}^{n}\frac{\alpha(G\boxtimes G^{i-1})}{\alpha(G^{i-1})}\leq\left[\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}\right]^{n}$ (2)
where $G^{0}$ is a trivial one-vertex graph, and $\sup$ is over all $W$. Letting
$\alpha^{*}(G)=\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}$ (3)
we deduce $\mathscr{C}(G)\leq\alpha^{*}(G)$. Hales in [8] showed that the supremum in (3) is a maximum and moreover, $\alpha^{*}(G)$ is equal to the fractional independence number of a graph $G$ (or Rosenfeld number), which is defined via a linear program as follows:
$\max\sum_{v\in\mathcal{V}(G)}w_{v}$ (4)
where the maximum is over all weights $w_{v}\geq 0$ such that $\sum_{v\in\mathcal{S}}w_{v}\leq 1$ for every clique $\mathcal{S}\subset\mathcal{V}(G)$. [^2] Some papers define the fractional independence number differently as the maximum sum of non-negative weights in $[0,1]$ that can be assigned to the vertices of a graph while ensuring that the total weight on any edge does not exceed 1. Here, we require the stronger condition that the total weight of any arbitrary clique does not exceed 1. In other words, $\alpha^{*}(G)$ is the fractional clique number of $G^{c}$ which is also equal to the fractional chromatic number of $G^{c}$. The fractional independence number is strictly greater than $\mathscr{C}(G)$ for many graphs $G$. Observe that the relaxation in equation (2) yields a weak bound in general because the left-hand side considers the expression ${\alpha(G\boxtimes W)}/{\alpha(W)}$ when $W=G^{i}$. In contrast, the right-hand side takes a maximum over all auxiliary graphs $W$.
There are also other upper bounds on $\mathscr{C}(G)$, such as Lovász number of a graph [9] or the Haemers number [10]. The Lovász number can be formulated as a semidefinite program and numerically approximated by the ellipsoid method in polynomial time in the number of vertices of $G$ [11]. The Haemers number of a graph $G$, $\mathcal{H}(G)$, is an upper bound on $\mathscr{C}(G)$ which considers the rank of particular matrices associated with the graph $G$ [10]. An upper bound on the Shannon capacity of a graph via a linear programming variation is given in [12]. In [13], a fractional version of Haemers bound is presented, and it is shown that this fractional version outperforms the Haemers bound. It is generally challenging to compute the exact values of Haemers number and the fractional Haemers number of a given graph $G$. However, these values are computed exactly [10, 14] for some special graphs, such as Johnson graphs with certain parameters. Please see [15, 16, 17, 6] for more related results and survey papers.
1.1 Our contribution
We consider two graphs, $G$ and $H$, and instead of writing individual upper bounds on the Shannon number of $G$ and $H$, we write a bound on the ratio of their capacities as follows:
$\frac{\mathscr{C}(G)}{\mathscr{C}(H)}$ $=\lim_{n\rightarrow\infty}\left(\frac{\alpha(G^{n})}{\alpha(H^{n})}\right)^{\frac{1}{n}}$ (5)
$=\lim_{n\rightarrow\infty}\left(\prod_{i=1}^{n}\frac{\alpha(G^{i}\boxtimes H^{n-i})}{\alpha(G^{i-1}\boxtimes H^{n-i+1})}\right)^{\frac{1}{n}}$
$=\lim_{n\rightarrow\infty}\left(\prod_{i=1}^{n}\frac{\alpha(G\boxtimes W_{i})}{\alpha(H\boxtimes W_{i})}\right)^{\frac{1}{n}}$ (6)
$\leq\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}$ (7)
where in (6) we set $W_{i}=G^{i-1}\boxtimes H^{n-i}$ and the supremum in (7) is taken over all graphs $W$ [^3] The expansion used here mimics the following one widely used in network information theory. Assuming $I(\cdot;\cdot)$ is Shannon’s mutual information, for arbitrarily distributed random variables $(M,Y_{1},Y_{2},\cdots,Y_{n},Z_{1},Z_{2},\cdots,Z_{n})$ we have $I(M;Y^{n})-I(M;Z^{n})=\sum_{i=1}^{n}\left(I(M;Y^{i},Z_{i+1}^{n})-I(M;Y^{i-1},Z_{i}^{n})\right)=\sum_{i=1}^{n}\left(I(M;Y_{i}|Y^{i-1},Z_{i+1}^{n})-I(M;Z_{i}|Y^{i-1},Z_{i+1}^{n})\right)$.
Given two graphs $G$, and $H$, we introduce the following notation for the term on the right-hand side of (7):
$\alpha^{*}(G|H)\triangleq\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)},$ (8)
where the supremum is over all graphs $W$. If $H$ is a one-vertex graph, $\alpha^{*}(G|H)$ is $\alpha^{*}(G)$, the fractional independence number of $G$. So, we call $\alpha^{*}(G|H)$ the relative fractional independence number of $G$ with respect to $H$.
After defining the notion of $\alpha^{*}(G|H)$, we show that the supremum in (8) is a maximum attained by a finite graph $W$. We give a computable characterization of $\alpha^{*}(G|H)$ in terms of a linear program. While the linear program generalizes the one in (4), its form is non-trivial and more involved than the one given in [8]. Moreover, for the special case of $G$ being a vertex-transitive graph, we show that $W=G^{c}$ is a maximizer in (8) and
$\alpha^{*}(G|H)=\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)}.$ (9)
If both $G$ and $H$ are vertex-transitive graphs, we can also write
$\alpha^{*}(G|H)=\frac{\chi_{f}(G^{c}\boxtimes H)}{|\mathcal{V}(H)|},$ (10)
where $\chi_{f}(\cdot)$ is the fractional coloring number. To the best of our knowledge, our characterization of $\alpha^{*}(G|H)$ for arbitrary graphs $G$ and $H$ leads to the first explicit non-trivial upper bound on the ratio of the zero-error capacity of two arbitrary graphs, as discussed later in Section 3.3. In particular, for a vertex-transitive graph $G$ and any arbitrary graph $H$ we obtain
$\frac{\mathscr{C}(G)}{\mathscr{C}(H)}\leq\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)}.$ (11)
To illustrate the above bound, we obtain explicit new bounds on the ratio of the zero-error Shannon capacity for two Cayley graphs (see Table 1). Next, we discuss the relation of $\alpha^{*}(G|H)$ with other invariants, such as the Lovász number of a graph. For a given graph $G$, using the fractional independence number of $G$ with respect to another graph $H$, we can give upper or lower bounds on the value of these invariants or compute some previously unknown invariants. As an application, we obtain lower bounds on the Shannon capacity of certain Johnson graphs, enabling us to determine the exact value of their Haemers number for the first time. Moreover, while the typical approach to proving lower bounds on graphs’ independence number (or the Shannon number) is by explicitly exhibiting an independent set, $\alpha^{*}(G|H)$ provides an indirect approach to proving such lower bounds. In particular, for any graph $H$, we have
$\alpha(H)\geq\frac{\alpha(G)}{\alpha^{*}(G|H)}$
which provides a non-constructive “lower bound” on the independence number of a graph without explicitly exhibiting the vertices of an independent set for $H$. This lower bound is useful if the structure of the graphs $G$ and $H$ can be utilized to compute $\alpha^{*}(G|H)$ from its linear programming representation.
Lastly, determining whether a homomorphism exists between two graphs is a fundamental problem in graph theory. The No-Homomorphism Lemma is a well-known tool in graph theory used to demonstrate the non-existence of a homomorphism between two graphs. Section 3.4 presents a stronger version of the No-Homomorphism lemma utilizing the relative fractional independence number. In some cases, our new lemma shows no homomorphism between two graphs, while the original version does not.
$\mathbf{G}$ $\mathbf{H}$ $\mathbf{\alpha^{*}(G|H)}$ Remarks
Theorem 9
$Cay(\mathbb{Z}_{n},\pm 1,\dots\pm k)$ $Cay(\mathbb{Z}_{m},\pm 1,\dots,\pm k)$ $n/m$ $1\leq 2k<n<m$
$m=\ell n+s(k+1)$
for integers $\ell,s\geq 0$
$Cay(\mathbb{Z}_{n},\pm 1,\cdots,\pm(2k+1))$ $Cay(\mathbb{Z}_{n},\pm 1,\cdots,\pm k)$ $1/2$ Example 1, $n\geq 4k+3$
$Cay(\mathbb{Z}_{n_{1}},\pm 1,\dots,\pm(r+1))$ $Cay(\mathbb{Z}_{n_{2}},\pm 1,\dots,\pm r)$ 1 Example 2
$n_{1}=(r+2)k+r$ $n_{2}=(r+1)k+r$ $r,k\in\mathbb{N}$
$C_{2k+1}$ $J(4k+2,3)$ $1/4$ Lemma 5
Appendix C, $n,m>1$
$C_{n}$ $C_{m}$ $n/(m-1)$ if $n$ even and $m$ odd
$n/m$ if $m$ even
$n/m$ if $n,m$ odd and $n\leq m$
$n/(m-1)$ if $n,m$ odd and $n>m$
Table 1: Values of $\alpha^{*}(G|H)$ for some graph pairs; $Cay$ is the Cayley graph, $C_{n}$ is a cycle of length $n$ and $J(n,k)$ is the Johnson graph.
1.2 Notation
Throughout this paper, we use capital letters such as $G,H$, and $W$ to denote graphs (finite, undirected graphs with no loops or multiple edges). The complement of a graph $G$, denoted by $G^{c}$, is a graph with the same vertices as in $G$, such that two distinct vertices of $G^{c}$ are adjacent iff they are not adjacent in $G$. We use $C_{k}$ to denote a cycle graph of length $k$. We show sets in calligraphic letters. We use the lowercase letters to either denote vertices of a graph (as in $u,v,v_{1},v_{2},x,y$) or real numbers (as in $w_{1},w_{2},...$). The bold letter $\mathbf{w}$ denotes a vector of real numbers. For graphs $G_{1}$ and $G_{2}$, $G=G_{1}+G_{2}$ is the disjoint union of $G_{1}$ and $G_{2}$, i.e., a graph whose vertex set is the union of vertex sets of $G_{1}$ and $G_{2}$ and whose edge set consists of edges which are in $G_{1}$ or in $G_{2}$.
We say that $G$ is an induced subgraph of $H$ if $G$ is formed by selecting a subset of the vertices from $H$ and then including all the edges from the parent graph $H$ that connect those chosen vertices. Given a graph $G$, let $\mathcal{I}(G)$ be the set of all independent sets of $G$. We also consider the empty set as a member of $\mathcal{I}(G)$. Finally, a homomorphism $\mathsf{g}:H\rightarrow G$ from a graph $H$ to a graph $G$ is a map $\mathsf{g}:\mathcal{V}(H)\rightarrow\mathcal{V}(G)$ such that $uv\in\mathcal{E}(H)$ implies $\mathsf{g}(u)\mathsf{g}(v)\in\mathcal{E}(G)$.
A graph is called a vertex-transitive (or node symmetric) graph if and only if for its any pair of nodes $v$ and $u$, there exists an automorphism [^4] An automorphism of a graph is a graph isomorphism with itself, i.e., a mapping from the vertices of the given graph $G$ back to vertices of $G$ such that the resulting graph is isomorphic with $G$. of the graph that maps $v$ to $u$ [18]. If $G$ and $H$ are vertex-transitive graphs, then so is $G^{c}$ and $G\boxtimes H$ (see [19, Lemma 1]).
We also need the definition of a Cayley graph:
Definition 1.
Let $G$ be an abelian group and $S$ be a subset of elements of $G$ such that $S=-S$ and $0\not\in S$. The Cayley graph $Cay(G,S)$ is a graph with the vertex set $G$ such that $a,b\in G$ are connected by an edge iff $a-b\in S$. Cayley graphs are basic examples of vertex-transitive graphs.
2 Characterizations and properties of $\alpha^{*}(G|H)$
This section discusses the characterizations and properties of $\alpha^{*}(G|H)$.
2.1 Characterizations of $\alpha^{*}(G|H)$
Definition 2.
We say that $\mathcal{S},\mathcal{T}\in\mathcal{I}(G)$ are disconnected in $G$ if $\mathcal{S}\cap\mathcal{T}=\emptyset$ and $\mathcal{S}\cup\mathcal{T}\in\mathcal{I}(G)$.
Definition 3.
Following [19], for a given graph $G$, the independence graph corresponding to $G$, denoted by $G^{*}$, is a simple graph whose vertices are elements of $\mathcal{I}(G)$ and two vertices are nonadjacent in $G^{*}$ iff their corresponding independent sets are disconnected. Equivalently, there is an edge between $\mathcal{S},\mathcal{T}\in\mathcal{I}(G)$ where $\mathcal{S}\neq\mathcal{T}$ iff $\mathcal{S}\cap\mathcal{T}\neq\emptyset$ or $\mathcal{S}\cup\mathcal{T}\notin\mathcal{I}(G)$. [^5] There is also a different definition of independence graph in the literature which we do not use here [20].
Definition 4.
Given two arbitrary graphs $H$ and $G$, let $\hat{H}$ be $(H^{*})^{c}$ with an added loop for the vertex corresponding to $\emptyset$ to itself. Then, we define $\mathcal{F}(H,G)$ as the class of all homomorphisms from $G^{c}$ to $\hat{H}$. Equivalently, $\mathcal{F}(H,G)$ is the set of functions $f:\mathcal{V}(G)\rightarrow\mathcal{I}(H)$ with the following property: $f(v_{1})$ and $f(v_{2})$ are disconnected in $H$ if there is no edge between $v_{1}$ and $v_{2}$ in $G$.
Remark 1.
If a homomorphism $\mathsf{g}$ from $H$ to $G$ exists, for any vertex $w\in\mathcal{V}(G)$, the set $\{v\in\mathcal{V}(H):\mathsf{g}(v)=w\}$ must be an independent set in $H$ (empty set is considered to be an independent set). Thus, the “inverse” of the homomorphism is a mapping $\mathsf{g}^{-1}:\mathcal{V}(G)\rightarrow\mathcal{I}(H)$ with the following properties:
  1. (i)
    $\mathsf{g}^{-1}(v_{1})$ and $\mathsf{g}^{-1}(v_{2})$ are disconnected in $H$ if there is no edge between $v_{1}$ and $v_{2}$ in $G$,
  2. (ii)
    The sets $\{\mathsf{g}^{-1}(v)\}$ for $v\in\mathcal{V}(G)$ form a partition of $\mathcal{V}(H)$.
If we relax the condition (ii) above, and just keep (i), we get the definition of $\mathcal{F}(H,G)$.
Now, we are ready to present our main theorem.
Theorem 1.
Assume that $G$ is a graph with $k$ vertices and $\mathcal{V}(G)=\{v_{1},v_{2},\cdots,v_{k}\}$. Then, we have
  1. (i)
    $\alpha^{*}(G|H)=\max_{\mathbf{w}}\sum_{i=1}^{k}w_{i},$ (12)
    where the maximization over $\mathbf{w}=(w_{1},w_{2},\cdots,w_{k})$ is subject to $w_{i}\geq 0$ and $\sum_{i=1}^{k}w_{i}|f(v_{i})|\leq 1$ for any $f\in\mathcal{F}(H,G)$. In other words, we require $\sum_{i=1}^{k}w_{i}|\mathcal{T}_{i}|\leq 1$ for any collection of sets $\mathcal{T}_{1},\cdots,\mathcal{T}_{k}\in\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected in $H$ if there is no edge between $v_{i}$ and $v_{j}$ in $G$.
  2. (ii)
    $\frac{1}{\alpha^{*}(G|H)}=\min_{w_{i}\geq 0,\sum_{i=1}^{k}w_{i}=1}\max_{f}\sum_{i=1}^{k}w_{i}|f(v_{i})|,$
    where the maximum is over all $f\in\mathcal{F}(H,G)$.
  3. (iii)
    Given a probability distribution $p(f)$ over $f\in\mathcal{F}(H,G)$, we can consider a random $F\in\mathcal{F}(H,G)$. We have
    $\alpha^{*}(G|H)=\min\max_{v\in\mathcal{V}(G)}\frac{1}{\mathbb{E}[|F(v)|]},$
    where the minimum is over all possible distributions on the set $\mathcal{F}(H,G)$.
  4. (iv)
    For every pair of graphs $G$ and $H$, there is some graph $W$ such that
    $\alpha^{*}(G|H)=\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$
    In other words, the supremum in the definition of $\alpha^{*}(G|H)$ is a maximum.
We give two different proofs for the above theorem. While the first proof given in Section 5.1 is longer, it provides insights into the form of the linear program. On the other hand, if we already know the form of the linear program, its correctness can be verified directly via the second proof presented in Appendix D (which is an extension of the proof presented in [8]).
Let us outline the first proof. We would like to maximize $\alpha(W\boxtimes G)/\alpha(W\boxtimes H)$. Take some arbitrary graph $W$. Let $\mathcal{B}$ be a maximum independent set for $W\boxtimes G$, i.e., $|\mathcal{B}|=\alpha(W\boxtimes G)$. Then, for every vertex $u$ of $W$, we form the set $\mathcal{B}_{u}=\{v\in\mathcal{V}(G):(u,v)\in\mathcal{B}\}$. Next, we partition the vertices of $W$ according to $\mathcal{B}_{u}$ (vertices with the same set $\mathcal{B}_{u}$ will be placed in the same class). Next, using this partition, we add some edges to the graph $W$ such that $\mathcal{B}$ is still a maximum independent set for $W\boxtimes G$, but $\alpha(W\boxtimes H)$ can potentially decrease. Once the edges are added, the graph $W$ will have a particular structure from which we can explicitly compute $\alpha(W\boxtimes H)$ and optimize over $W$. In Appendix A, we compute the $\alpha^{*}(C_{7}|C^{c}_{7})$ as a warm-up example.
2.2 Properties of $\alpha^{*}(G|H)$
From the definition of $\alpha^{*}(G|H)$ in (8), some immediate observations can be made: for any graph $G$, we have $\alpha^{*}(G|H)=\alpha^{*}(G)$ if $H$ is a complete graph. Also, $\alpha^{*}(G|G)=1$. As another example, assume that the graph $H$ is universal [21], i.e., $\alpha(H\boxtimes W)=\alpha(H)\alpha(W)$ for all $W$ [^6] Examples of universal graphs are perfect graphs. A graph is a perfect graph iff its induced subgraphs include neither an odd cycle of length greater than five nor an odd anti-cycle of length greater than five. Examples of perfect graphs are complete graphs or a cycle of even length, $C_{2k}$. For a universal graph $H$, we have
$\alpha^{*}(G|H)=\frac{1}{\alpha(H)}\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}=\frac{\alpha^{*}(G)}{\alpha(H)}.$
Also, if $G=G_{1}\boxtimes G_{2}$ where $G_{1}$ is universal, then $\alpha^{*}(G|H)=\alpha(G_{1})\alpha^{*}(G_{2}|H)$. And if $H=H_{1}\boxtimes H_{2}$ where $H_{1}$ is universal, then $\alpha^{*}(G|H)=\frac{\alpha^{*}(G|H_{2})}{\alpha(H_{1})}$. Some properties of $\alpha^{*}(G|H)$ are given in the following theorem.
Theorem 2.
For any graphs $G_{1},G_{2},H_{1}$ and $H_{2}$ we have:
  1. (i)
    $\alpha^{*}(G_{1}\boxtimes G_{2}|H_{1}\boxtimes H_{2})\leq\alpha^{*}(G_{1}|H_{1})\alpha^{*}(G_{2}|H_{2})$.
  2. (ii)
    $\frac{1}{\alpha^{*}(H|G)}\leq\frac{\alpha^{*}(G|W)}{\alpha^{*}(H|W)}\leq\alpha^{*}(G|H)$ for any arbitrary graph $W$.
  3. (iii)
    $\alpha^{*}(G_{1}+G_{2}|H)\leq\alpha^{*}(G_{1}|H)+\alpha^{*}(G_{2}|H)$.
  4. (iv)
    $\alpha^{*}((G_{1}+G_{2})^{c}|H)=\max(\alpha^{*}(G_{1}^{c}|H),\alpha^{*}(G_{2}^{c}|H)).$
  5. (v)
    $\frac{1}{\alpha^{*}(G|H_{1})}+\frac{1}{\alpha^{*}(G|H_{2})}\leq\frac{1}{\alpha^{*}(G|H_{1}+H_{2})}$. Moreover, equality holds if $G$ is a vertex-transitive graph.
Remark 2.
The first part of Theorem 2 implies that for any three graphs $G$, $H$ and $T$, we have $\alpha^{*}(G\boxtimes T|H\boxtimes T)\leq\alpha^{*}(G|H)$. Next, note that the first part of Theorem 2 holds with equality if $H_{1}$ and $H_{2}$ are single vertices. However, strict inequality may hold in the first part of Theorem 2 if $G_{1}$ and $G_{2}$ are single vertices as
$\alpha^{*}(G_{1}\boxtimes G_{2}|H_{1}\boxtimes H_{2})<\alpha^{*}(G_{1}|H_{1})\alpha^{*}(G_{2}|H_{2})$
is equivalent with
$\alpha(H_{1}\boxtimes H_{2})>\alpha(H_{1})\alpha(H_{2})$
which may occur.
A proof of the above theorem is given in Section 5.2. To continue, we need a property of the fractional independence number given in the following lemma.
Lemma 1.
For any arbitrary graphs $G_{1},G_{2},G_{2},\cdots,G_{r}$, there is some graph $W$ such that
$\alpha^{*}(G_{i})$ $=\frac{\alpha(G_{i}\boxtimes W)}{\alpha(W)},\qquad\forall i.$
That is, a common maximizer $W$ for the optimization problems involving $\alpha^{*}(G_{i})$ exists.
Proof of the above lemma is given in Section 5.3.
Corollary 1.
Given two arbitrary graphs $G$ and $H$, Lemma 1 implies the existence of some $W$ such that
$\frac{\alpha^{*}(G)}{\alpha^{*}(H)}=\frac{\frac{\alpha(G\boxtimes W)}{\alpha(W)}}{\frac{\alpha(H\boxtimes W)}{\alpha(W)}}=\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$
Thus, from the definition of $\alpha^{*}(G|H)$ we obtain
$\alpha^{*}(G|H)\geq\frac{\alpha^{*}(G)}{\alpha^{*}(H)}.$ (13)
Also, we have:
$\alpha^{*}(G\boxtimes H|H)=\alpha^{*}(G).$ (14)
Proof.
One direction follows from (13). The other direction follows from
$\alpha^{*}(G\boxtimes H|H)$ $=\max_{W}\frac{\alpha(G\boxtimes(H\boxtimes W))}{\alpha(H\boxtimes W)}\leq\max_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}=\alpha^{*}(G)$
where the inequality follows from the fact that we are relaxing the set of graphs of the form $H\boxtimes W$ to all graphs $W$. ∎
The inequality (13) motivates us to ask whether $\alpha^{*}(G|H)$ serves as an upper bound on the ratio of other graph invariants for $G$ and $H$. The following theorem addresses this question.
Theorem 3.
For any graphs $G$ and $H$,
$\alpha^{*}(G|H)\geq\frac{X(G)}{X(H)}\geq\frac{1}{\alpha^{*}(H|G)}$
where $X(G)$ can be the independence number of $G$, the fractional independence number of $G$, the Lovász number of $G$ [9], Schrijver’s variant of the Lovász number of $G$ [22, 23] or Szegedy’s variant of the Lovász number of $G$ [24].
A proof of the above theorem is given in Section 5.4. Theorem 3 does not hold when $X(G)$ is the Haemers number of a graph $G$, $\mathcal{H}(G)$ (in Appendix B, we give two graphs $G$ and $H$, such that $\frac{\mathcal{H}(G)}{\mathcal{H}(H)}>\alpha^{*}(G|H)$). We do not know if Theorem 3 holds when $X(G)$ is the fractional Haemers number of a graph $G$.
Corollary 2.
Using Remark 2, we conclude the following chain of inequalities for Shannon capacity:
$\alpha^{*}(G|H)\geq\sup_{T}\frac{\mathscr{C}(G\boxtimes T)}{\mathscr{C}(H\boxtimes T)}\geq\frac{\mathscr{C}(G)}{\mathscr{C}(H)}\geq\inf_{T}\frac{\mathscr{C}(G\boxtimes T)}{\mathscr{C}(H\boxtimes T)}\geq\frac{1}{\alpha^{*}(H|G)}.$
Observe that for any two graphs $G$ and $H$ we have $\mathscr{C}(G\boxtimes T)\geq\mathscr{C}(G)\mathscr{C}(T)$ and the strict inequality may occur [25]. We do not know if the supremum over $T$ in
$\sup_{T}\frac{\mathscr{C}(G\boxtimes T)}{\mathscr{C}(H\boxtimes T)}$
may occur for some non-trivial graph $T$, and if so, for which graph pairs $G$ and $H$.
2.3 Computation of $\alpha^{*}(G|H)$ for a vertex-transitive graph $G$
The following result is well-known (e.g., see [26, Lemma 6]):
Lemma 2.
For any graph $G$,
$\alpha^{*}(G)=\chi_{f}(G^{c})\geq\frac{|\mathcal{V}(G)|}{\alpha(G^{c})}$
where $\chi_{f}(G)$ is the fractional chromatic number of $G$. Moreover, equality holds for a vertex-transitive $G$. In other words, for a vertex-transitive graph $G$, $W=G^{c}$ is a maximizer of
$\max_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}$
as $\alpha(G\boxtimes G^{c})=|\mathcal{V}(G)|$ for a vertex-transitive graph $G$ [19, Corollary 1(b)].
Observe that for any arbitrary graph $G$, using the special choice of $W=G^{c}$ we obtain the lower bound
$\alpha^{*}(G|H)=\max_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}\geq\frac{\alpha(G\boxtimes G^{c})}{\alpha(H\boxtimes G^{c})}\geq\frac{|\mathcal{V}(G)|}{\alpha(H\boxtimes G^{c})}$ (15)
where $\alpha(G^{c}\boxtimes G)\geq|\mathcal{V}(G)|$ follows from the fact that the set $\{(v,v):v\in\mathcal{G}\}$ is an independent set in $G^{c}\boxtimes G$. We provide an extension of the Lemma 2 for a vertex-transitive graph $G$ as follows:
Theorem 4.
For any graphs $G$ and $H$, where $\mathcal{V}(G)=\{v_{1},\dots,v_{k}\}$, we have
$\alpha^{*}(G|H)\geq\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)},$ (16)
Moreover, the lower bound in (16) is tight (holds with equality) if $G$ is vertex-transitive. In other words, for a vertex-transitive $G$, the graph $W=G^{c}$ is a maximizer
$\alpha^{*}(G|H)=\max_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$
Furthermore, for any graph $G$ on $k$ vertices $\mathcal{V}(G)=\{v_{1},\dots,v_{k}\}$, we have $\alpha(G^{c}\boxtimes H)=\max(\sum_{i=1}^{k}|\mathcal{T}_{i}|)$ where the maximum is over all collection of sets $\mathcal{T}_{1},\cdots,\mathcal{T}_{k}\in\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected in $H$ if there is no edge between $v_{i}$ and $v_{j}$ in $G$.
A proof of the above theorem is given in Section 5.6.
Corollary 3.
If $G$ is vertex-transitive, we have
$\alpha^{*}(G|H)=\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)}=\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}\cdot\frac{|\mathcal{V}(H)|}{\alpha(G^{c}\boxtimes H)}\leq\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}\cdot\alpha^{*}(H^{c}|G^{c}).$ (17)
Equality holds in (17) if both $G$ and $H$ are vertex-transitive. In Theorem 9 we compute $\alpha^{*}(G|H)$ for certain Cayley graphs that are vertex-transitive. The above observation implies that we can also obtain the relative fractional independence number for their compliments. If both $G$ and $H$ are vertex-transitive graphs, $G^{c}\boxtimes H$ will also be vertex-transitive and we can write
$\alpha^{*}(G|H)=\frac{|V(G)|}{\alpha(G^{c}\boxtimes H)}=\frac{\chi_{f}(G^{c}\boxtimes H)}{|\mathcal{V}(H)|}$
where we used Theorem 4 and Lemma 2.
Corollary 4.
If both $G$ and $H$ are vertex-transitive, we have
$\alpha^{*}(G\star H)\geq\alpha(G\boxtimes H)$
where $G\star H$ is the disjunctive product (also known as the co-normal product or OR product) between two graphs $G$ and $H$, defined as follows: we connect $(u,v)\in\mathcal{V}(G)\times\mathcal{V}(H)$ and $(u^{\prime},v^{\prime})\in\mathcal{V}(G)\times\mathcal{V}(H)$ if $uu^{\prime}\in\mathcal{E}(G)$ or $vv^{\prime}\in\mathcal{E}(H)$. This inequality may be new as we have not seen it reported elsewhere.
Proof.
Consider the inequality $\alpha^{*}(G^{c}|H)\alpha^{*}(H|G^{c})\geq 1$ which implies that
$\alpha(G\boxtimes H)\alpha(H^{c}\boxtimes G^{c})\leq|\mathcal{V}(G)||\mathcal{V}(H)|.$
Since
$\alpha^{*}((G^{c}\boxtimes H^{c})^{c})=\frac{|\mathcal{V}(G)||\mathcal{V}(H)|}{\alpha(G^{c}\boxtimes H^{c})}$
we deduce that
$\alpha^{*}((G^{c}\boxtimes H^{c})^{c})\geq\alpha(G\boxtimes H)$
Observe that $(G^{c}\boxtimes H^{c})^{c}=G\star H$. This completes the proof.
Observe that $G\boxtimes H$ is a subgraph of $G\star H$. Thus, the inequality in the above corollary immediately implies the following chain of inequalities for arbitrary vertex-transitive graphs $G$ and $H$:
$\alpha^{*}(G)\alpha^{*}(H)=\alpha^{*}(G\boxtimes H)\geq\alpha^{*}(G\star H)\geq\alpha(G\boxtimes H)\geq\alpha(G\star H)=\alpha(G)\alpha(H).$
Remark 3.
Theorem 4 allows us to use known results to obtain $\alpha^{*}(G|H)$ for certain special graphs. For instance, it is known [19, Theorem 1][8] that for $j\geq k$ we have $\alpha(C_{2j+1}\boxtimes C_{2k+1})=jk+\lfloor\frac{k}{2}\rfloor$ where $C_{k}$ is a cycle graph of size $k$. Since a cycle graph is vertex-transitive, Theorem 4 allows us to compute the relative independence number between the complement of an odd cycle with respect to another odd cycle. More results on the independence of the strong product of three odd cycles are reported in [27] and can be used to compute the relative fractional independence numbers for other pairs of graphs. Similar results are available for other graphs. For instance, it is proved in [28] that
$\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)\boxtimes Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k))=\left\lfloor\frac{\left\lfloor\frac{n}{k+1}\right\rfloor n}{k+1}\right\rfloor.$
We also remark that some of the known results in the literature follow immediately from our results. For instance, Lemma 2 in [29] states that $2\alpha(G\boxtimes C_{2k+1})\leq(2k+1)\alpha(G)$ where $C_{2k+1}$ is a cycle of size $12k+$ 1. Our proof of this fact is as follows:
$\frac{\alpha(C^{c}_{2k+1})}{\alpha(G)}\leq\alpha^{*}(C^{c}_{2k+1}|G)=\frac{2k+1}{\alpha(C_{2k+1}\boxtimes G)}.$
As another example, Theorem 6 in [29] states that $\chi(P_{5}\boxtimes C_{2k+1})>5$ where $P_{5}$ is the Petersen graph and where $\chi(\cdot)$ is the coloring number of a graph. Our proof of this fact is as follows:
$\frac{\vartheta(C^{c}_{2k+1})}{\vartheta(P_{5})}\leq\alpha^{*}(C^{c}_{2k+1}|P_{5})=\frac{\chi_{f}(P_{5}\boxtimes C_{2k+1})}{10}\leq\frac{\chi(P_{5}\boxtimes C_{2k+1})}{10}.$
where $\vartheta(G)$ is the Lovász number of $G$. Note that $\vartheta(C^{c}_{2k+1})>2$ since using the fact that the cycle graph is vertex-transitive we have
$\vartheta(C^{c}_{2k+1})=\frac{2k+1}{\vartheta(C_{2k+1})}>2.$
Next, note that $\vartheta(P_{5})=\alpha(P_{5})=4$ since $P_{5}$ is a Kneser graph. Thus, we obtain the desired inequality. As another example, Proposition 2.3 in [30] states that for positive integers $k,n,p$ such that $k<n$ we have
$\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{\boxtimes p})\leq\left\lfloor\frac{n}{k+1}\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{\boxtimes(p-1)})\right\rfloor$
where $G^{\boxtimes p}$ is the strong product of $G$ with itself $p$ times. Our proof of this fact is as follows:
$\frac{\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{c})}{\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{\boxtimes(p-1)})}$ $\leq\alpha^{*}(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{c}|Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{\boxtimes(p-1)})$
$=\frac{n}{\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)\boxtimes Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{\boxtimes(p-1)})}.$
The desired inequality follows from $\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)^{c})\geq k+1$.
Remark 4.
Assume that $G$ is a single vertex, and $H$ is an arbitrary graph. By Theorem 4, we have
$\alpha^{*}(G|H)=\frac{1}{\alpha(H)}$
So, in this case, computing $\alpha^{*}(G|H)$ is equivalent to computing $\alpha(H)$. Therefore, since computing $\alpha(H)$ is an NP-hard problem, so is computing $\alpha^{*}(G|H)$. Additionally, because of the hardness of approximation of $\alpha(H)$, no efficient upper bound for $\alpha^{*}(G|H)$ can be found in the general case.
We proceed with another application of Theorem 4. Note that for any graph $G$, the independence graph $G^{*}$ is defined in Definition 3. For any vertex $v\in G^{*}$, we can assign a number $N_{v}$ which is the size of the independent set in $G$ associated with the vertex $v\in G^{*}$.
Theorem 5.
For any graph $G$, if $H$ is an induced subgraph of $G^{*}$ then
$\alpha^{*}(H|G)\leq\left(\min_{v\in\mathcal{V}(H)}N_{v}\right)^{-1}.$
Moreover, if $H$ is an induced subgraph of $G^{*}$ that is also vertex-transitive, then
$\alpha^{*}(H|G)\leq\frac{|\mathcal{V}(H)|}{\sum_{v\in\mathcal{V}(H)}N_{v}}.$
Example 1.
Let $G=Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm k)$ for some $n\geq 4k+3$. Consider independent sets of the form $S_{i}=\{i,i+k+1\}$ for $1\leq i\leq n$. The induced subgraph in $G^{*}$ on these independent sets is $H=Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm(2k+1))$, yielding
$\alpha^{*}\big{(}Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm(2k+1))|Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm k)\big{)}\leq\frac{1}{2}.$
On the other hand,
$\alpha^{*}\big{(}Cay(\mathbb{Z}_{n},\pm 1,\cdots,\pm(2k+1))|Cay(\mathbb{Z}_{n},\pm 1,\cdots,\pm k)\big{)}\geq\frac{\alpha^{*}(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm(2k+1))}{\alpha^{*}(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm k))}=\frac{1}{2}.$
Thus,
$\alpha^{*}\big{(}Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm(2k+1))|Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\cdots,\pm k)\big{)}=\frac{1}{2}.$
Theorem 5 immediately follows from Theorem 4 and the following result of [19]:
Theorem 6 (Corrected form of Lemma 4 in [19]).
For any graphs $G$ and $W$ we have
$\alpha(G\boxtimes W)$ $=\max_{H\subset G^{*}}\big{\{}\alpha(H\boxtimes W)\min_{v\in\mathcal{V}(H)}N_{v}\big{\}}$ (18)
where $H\subset G^{*}$ means that $H$ is an induced subgraph of $G^{*}$. Next, if $H$ is an induced subgraph of $G^{*}$, then
$\alpha(G\boxtimes H^{c})\geq\sum_{v\in\mathcal{V}(H)}N_{v}$
Finally, a method for the construction of non-vertex-transitive graphs $T$ and $H$ for which
$\alpha^{*}(T|H)=\frac{|\mathcal{V}(T)|}{\alpha(T^{c}\boxtimes H)}$
holds is given in the following proposition.
Proposition 1.
Take an arbitrary vertex-transitive graph $G$ and an arbitrary graph $H$. Let $T$ be obtained by adding some new edges to $G$. The addition of new edges to graph $G$ is done in such a way that $\alpha(T^{c}\boxtimes H)=\alpha(G^{c}\boxtimes H)$ then
$\alpha^{*}(T|H)=\alpha^{*}(G|H)=\frac{|\mathcal{V}(T)|}{\alpha(T^{c}\boxtimes H)}$
Proof.
Since $T\in Expand(G)$, from Lemma 3 we have $\alpha^{*}(G|H)\geq\alpha^{*}(T|H)$. To show the reverse direction, note that
$\alpha^{*}(G|H)=\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)}=\frac{|\mathcal{V}(G)|}{\alpha(T^{c}\boxtimes H)}\leq\alpha^{*}(T|H).$
This shows $\alpha^{*}(G|H)=\alpha^{*}(T|H)$. ∎
2.3.1 The expand operations
Finally, we provide another definition that is useful in studying $\alpha^{*}(G|H)$.
Definition 5 (Expand Operations).
Let $G$ be an arbitrary graph. We define $Expand(G)$ as the set of all graphs that can be obtained by a sequence of the following three operations on $G$:
  1. (I)
    Remove a vertex $v$ from $G$.
  2. (II)
    Replace a vertex $v$ of $G$ by a clique of arbitrary size. In other words, put a clique of size $k$ instead of vertex $v$ and connect these $k$ new vertices to all neighbors of $v$.
  3. (III)
    Add a new edge to $G$.
Remark 5.
Note that by the above operations, one can merge (contract) any two vertices $v$ and $u$ in $G$. To achieve this, we only need to delete one of the vertices, for instance $v$, by applying an operation of type (I). Then, we add new edges between $u$ and the neighbors of $v$ by applying an operation of type (III). This allows us to merge the two vertices $v$ and $u$ into a single vertex.
One can easily verify that all the above operations, when applied to a graph $G$, will not increase $\alpha(G\boxtimes W)$ for any choice of graph $W$. Consequently, we obtain the following lemma:
Lemma 3.
If $G^{\prime}\in Expand(G)$ and $H^{\prime}\in Expand(H)$, then we have
$\alpha^{*}(G^{\prime}|H)\leq\alpha^{*}(G|H)\leq\alpha^{*}(G|H^{\prime}).$
Corollary 5.
If $G\in Expand(H)$, then $\alpha^{*}(G|H)\leq\alpha^{*}(H|H)=1$.
Corollary 6.
If $G\in Expand(H)$ and $\alpha(G)\geq\alpha(H)$, then $\alpha^{*}(G|H)=1$.
Proof.
Since $G\in Expand(H)$, we have $\alpha^{*}(G|H)\leq 1$. On the other hand, $\alpha^{*}(G|H)\geq\frac{\alpha(G)}{\alpha(H)}\geq 1$.
Example 2.
As an example, we compute the relative fractional independence number of a certain Cayley graph with respect to another Cayley graph. Let $G=Cay(\mathbb{Z}_{(r+2)k+r},\pm 1,\pm 2,\dots,\pm(r+1))$ and $H=Cay(\mathbb{Z}_{(r+1)k+r},\pm 1,\pm 2,\dots,\pm r)$ for natural numbers $r$ and $k$. We show that $\alpha^{*}(G|H)=1.$
Note that
$\alpha(G)=\left\lfloor\frac{(r+2)k+r}{r+2}\right\rfloor=k,\qquad\alpha(H)=\left\lfloor\frac{(r+1)k+r}{r+1}\right\rfloor=k.$
To utilize Corollary 6, we need to show that we can derive $G$ from $H$ by using the expand operations. A demonstration for the special case of $k=3,r=1$ is given in Figure 1. We proceed as follows: first label the $i$-th vertex of $H$ with
$u_{i+\left\lceil\frac{i}{r+1}\right\rceil-1}$
for $i=1,2,\cdots,(r+1)k+r$. In other words, the sequences of vertex labels are
$u_{1},u_{2},\cdots,u_{r+1},u_{r+3},u_{r+4},\cdots,u_{2r+3},u_{2r+5},u_{2r+6},...$
with no vertex labeled as $u_{(r+2)i}$. Now, for any $1\leq i\leq k$, replace each vertex $u_{(r+2)i-1}$ with a clique of size two and label the two new vertices with $u_{(r+2)i-1}$ and $u_{(r+2)i}$. The resulting graph has $(r+2)k+r$ vertices. The neighbors of each vertex is as follows:
$N(u_{i})=\begin{cases}\{u_{i-(r+1)},\dots,u_{i-1},u_{i+1},\dots,u_{i+r}\},&\text{if $i$ is divisible by $(r+2)$},\\ \{u_{i-(r+1)},\dots,u_{i-1},u_{i+1},\dots,u_{i+r+1}\},&\text{if the remainder of $i$ divided by $(r+2)$ is between $1$ and $r$, }\\ \{u_{i-r},\dots,u_{i-1},u_{i+1},\dots,u_{i+r+1}\},&\text{if the remainder of $i$ divided by $(r+2)$ is $r+1$}.\end{cases}$
This is a spanning subgraph of $G$. Therefore, by adding the missing edges to this graph, one can obtain $G=Cay(\mathbb{Z}_{(r+2)k+r},\pm 1,\pm 2,\dots,\pm(r+1))$.
Figure 1: Replacing vertices 2, 5, and 8 with a clique of size two, and then adding extra edges (1,9), (2,10), (3,5), and (6,8) yields the graph $Cay(\mathbb{Z}_{10},\pm 1,\pm 2)$.
We continue by a remark about Corollary 6:
Remark 6.
Take two arbitrary graphs, $G$ and $H$ and two natural numbers $p$ and $q$. Let
$H^{\prime}=\underbrace{H+H+\cdots+H}_{p\text{ times}}\,\,\text{and}\,\,G^{\prime}=\underbrace{G+G+\cdots+G}_{q\text{ times}}.$
Then, for any graph $W$, $\alpha(H^{\prime}\boxtimes W)=p\cdot\alpha(H\boxtimes W)$ and $\alpha(G^{\prime}\boxtimes W)=q\cdot\alpha(G\boxtimes W)$, and therefore $\alpha(G|H)\leq\frac{p}{q}$ is equivalent to $\alpha(G^{\prime}|H^{\prime})\leq 1$. Therefore, one may potentially utilize Corollary 6 to prove $\alpha(G|H)=\frac{p}{q}$.
Next, we have the following result:
Theorem 7.
Take arbitrary graphs $G$ and $H$, and let $\mathcal{S}$ be an independent set of $G^{c}\boxtimes H$. Define $\mathcal{S}_{G}\subset\mathcal{V}(G)$ as the set of vertices of $G$ used in the independent set as follows:
$\mathcal{S}_{G}=\{u\in\mathcal{V}(G):(u,v)\in\mathcal{S}\text{ for some }v\in\mathcal{V}(H)\}.$
Let $G^{\prime}$ be the induced subgraph of $G$ on vertices $\mathcal{S}_{G}$. Similarly, we define $H^{\prime}$ as the induced graph on the vertices of $H$ used in the independent set $\mathcal{S}$. Then, $G^{\prime}\in Expand(H^{\prime})$ and consequently, $\alpha^{*}(G^{\prime}|H^{\prime})\leq 1$. In particular, since $H^{\prime}\in Expand(H)$, we also obtain $G^{\prime}\in Expand(H)$ and $\alpha^{*}(G^{\prime}|H)\leq 1$.
A proof of the above theorem is given in Section 5.5.
2.3.2 $G\in Expand(H)$ vs $\alpha^{*}(G|H)\leq 1$
From Lemma 3, it follows that if $G\in Expand(H)$, then $\alpha^{*}(G|H)\leq\alpha^{*}(H|H)=1$. One might ask if the converse is also true, i.e., whether $\alpha^{*}(G|H)\leq 1$ implies that $G\in Expand(H)$. We show that the converse statement is false in general, but holds for a certain class of Cayley graphs (which includes all cycles) and for all perfect graphs.
Theorem 8.
In the following cases $\alpha^{*}(G|H)\leq 1$ implies $G\in Expand(H)$:
  1. i.
    If $G$ is the Cayley graph $Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)$, where $n=c(k+1)+1$ or $n=c(k+1)$ for an arbitrary integer $c$, and $H$ is an arbitrary graph
  2. ii.
    If $G$ is a perfect graph and $H$ is an arbitrary graph.
A proof of the above theorem is given in Section 5.7. Note that by setting $k=1$ in the first part of Theorem 8, we observe that Theorem 8 holds when $G$ is a cycle.
Next, we construct an example for which $\alpha^{*}(G|H)\leq 1$ but $G\notin Expand(H)$. Set $G=Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\pm 3)$ and $H=C_{n}$ where $n$ is an arbitrary odd number greater than $7$. Then, Example 1 yields $\alpha^{*}(G|C_{n})=\frac{1}{2}$ which is equivalent to
$\alpha^{*}(G+G|C_{n})=1.$
On the other hand, $G+G\notin Expand(C_{n})$ follows from Lemma 4 below as for any odd $n$ we have
$\alpha^{*}(G+G)=2\alpha^{*}(G)=\frac{n}{2}>\left\lfloor\frac{n}{2}\right\rfloor.$
Lemma 4.
If $T$ is a disconnected graph and $T\in Expand(C_{n})$, then
$\alpha^{*}(T)\leq\alpha(C_{n})=\left\lfloor\frac{n}{2}\right\rfloor$
Proof.
Definition 5 gives three types of operations. One can argue if $T=Expand(H)$, then one can also reach $T$ from $H$ by first performing some operations of type (I), then some operations of type (II), and finally some operations of type (III). To see this, observe that the operations of adding an edge (type (III)) can always be postponed to the very end. Then, if we have a sequence of operations of type (I) or (II), the operations of removing a vertex (type (I)) can all be performed first at the beginning.
Now, observe that since $T$ is a disconnected graph while $C_{n}$ is a connected graph, we must utilize the vertex deletion operation (operation of type (I)) in the first stage. This is because the second and third Expand operations alone cannot decrease the connectivity of a graph. Let $H^{\prime}$ be the graph obtained from removing at least one vertex from $H=C_{n}$. Then $T$ will be in the Expand of $H^{\prime}$, and therefore $\alpha^{*}(T|H^{\prime})\leq 1$. Note, however, that the graph $H^{\prime}$ will be a perfect graph (and thus a universal graph); removal of any one vertex from $C_{n}$ will make it a perfect graph. Therefore, $\alpha^{*}(T|H^{\prime})=\alpha^{*}(T)/\alpha(H^{\prime})$. Thus,
$\alpha^{*}(T)\leq\alpha(H^{\prime})\leq\alpha(C_{n})=\left\lfloor\frac{n}{2}\right\rfloor.$
Finding necessary or sufficient conditions for $G=Expand(H)$ is an interesting open problem. As an example, we give such a condition below. Let $\omega(G)=\alpha(G^{c})$ be the clique number of $G$ (size of the maximum clique in $G$).
Proposition 2.
If $G\in Expand(H)$, then one can assign a non-negative integer $s(v)$ to every vertex in $H$ such that
$|\mathcal{V}(G)|=\sum_{v\in\mathcal{V}(H)}s(v).$
Moreover, for every clique $C$ in $H$ we have
$\sum_{v\in C}s(v)\leq\omega(G).$
Proof.
Note that operation (III) in the expand operations can only increase the size of the maximum clique. Thus, after performing operations of type (I) and (II) on $H$, we cannot reach a graph whose clique number is larger $\omega(G)$. Assume that in operations (I) and (II) we replace every vertex $v$ of a graph $H$ with $s(v)$ copies ($s(v)=0$ is equivalent to removing the vertex, and $s(v)=1$ is equivalent with keeping the vertex untouched while $s(v)>1$ is equivalent to replicating the vertex). Then, for every clique $C$ in $H$ we must have
$\sum_{v\in C}s(v)\leq\omega(G).$
On the other hand, we must have
$|\mathcal{V}(G)|=\sum_{v\in\mathcal{V}(H)}s(v)$
as $\sum_{v\in\mathcal{V}(H)}s(v)$ is the total number of vertices after performing operations of type (I) and (II). ∎
3 Applications of the relative fractional independence number
In this section, we explore some applications of $\alpha^{*}(G|H)$.
3.1 New lower bounds on the Shannon capacity of certain Johnson graphs and the exact value of their Haemers number
Let $J(n,3)$ be the Johnson graph whose vertices are the $3$-subsets of $\{1,\dots,n\}$, and two vertices are adjacent if their intersection has one element. [^7] We use [10]’s definition Some papers call this graph’s complement the Johnson graph. Haemers [10] showed that $\mathcal{H}(J(n,3))\leq n$ and the equality holds for $n=4k$. Using the relative fractional independence number, we compute a new lower bound for $\mathscr{C}(J(n,3))$ for $n=14,18,22,26$. This new lower bound implies that $\mathcal{H}(J(n,3))=n$ for $n=14,18,22,26$, which is unknown in the literature. [^8] For $n=4k$, if we partition the underlying $n$-set into classes of size four, then all $3$-subsets, which are subsets of one of these classes, form an independent set of size $n$ in $J(n,3)$. Hence, $\alpha(J(n,3))\geq n$ when $n$ is divisible by $4$. Since $\mathcal{H}(J(n,3))\leq n$, [10] deduces that $\alpha(J(n,3))=\mathscr{C}(J(n,3))=\mathcal{H}(J(n,3))$. However, when $n$ is not divisible by $4$, the exact value of $\mathcal{H}(J(n,3))$ was unknown in [10].
We have $\frac{\mathscr{C}(C_{2k+1})}{\mathscr{C}(J(4k+2,3))}\leq\alpha^{*}(C_{2k+1}|J(4k+2,3))$. We first prove the following result:
Lemma 5.
We have
$\alpha^{*}(C_{2k+1}|J(4k+2,3))=\frac{1}{4}.$ (19)
Proof.
We know that
$\frac{\alpha(C_{2k+1})}{\alpha(J(4k+2,3))}\leq\alpha^{*}(C_{2k+1}|J(4k+2,3)).$
Note that $\alpha(C_{2k+1})=k$ and $\alpha(J(4k+2,3))=4k$ [31]. So, we obtain $\frac{1}{4}\leq\alpha^{*}(C_{2k+1}|J(4k+2,3)).$ Next, we have
$\alpha^{*}(C_{2k+1}|J(4k+2,3))=\frac{2k+1}{\max\sum^{2k+1}_{i=1}|\mathcal{T}_{i}|}.$
Now we give a valid assignment of the independent sets of $J(4k+2,3)$ to the vertices of $C_{2k+1}$. We assign the set $\mathcal{T}^{*}_{i}$ to $v_{i}$ for $11\leq i\leq 2k+$ 1 as follows: for $i\neq 2k+1$, let $\mathcal{T}^{*}_{i}=$ all the $3$-subsets of $\{2i-1,2i,2i+1,2i+2\}$ and $\mathcal{T}^{*}_{2k+1}=$ all the $3$-subsets of $\{4k+1,4k+2,1,2\}$. Then, we have
$\alpha^{*}(C_{2k+1}|J(4k+2,3))=\frac{2k+1}{\max\sum^{2k+1}_{i=1}|\mathcal{T}_{i}|}\leq\frac{2k+1}{\sum^{2k+1}_{i=1}|\mathcal{T}^{*}_{i}|}=\frac{2k+1}{4(2k+1)}=\frac{1}{4}.$
Equation (19) implies $\mathscr{C}(C_{2k+1})\leq\frac{1}{4}\mathscr{C}(J(4k+2,3))\leq\frac{1}{4}\mathcal{H}(J(4k+2,3))$, yielding the lower bounds $\mathscr{C}(J(4k+2,3))\geq 4\mathscr{C}(C_{2k+1})$ and $\mathcal{H}(J(4k+2,3))\geq\lceil 4\mathscr{C}(C_{2k+1})\rceil$ (we take the ceiling as $\mathcal{H}(J(n,3))$ is an integer number). Since $\mathscr{C}(C_{7})\geq 3.2578$ [32], $\mathscr{C}(C_{9})\geq 4.32$ [33], $\mathscr{C}(C_{11})\geq 5.2895$ [33], $\mathscr{C}(C_{13})\geq 6.2743$ [34], we get new lower bounds on the Shannon capacity and can conclude $\mathcal{H}(J(n,3))=n$ for $n=14,18,22,26$. Note that we obtained a lower bound on $\mathscr{C}(J(4k+2,3)$ indirectly, without showing an independent set for this graph (or its powers). A better lower bound for $\mathscr{C}(C_{2k+1})$ would also improve the lower bound of $\mathscr{C}(J(4k+2,3)$. Also, Theorem 1.1 in [3] states that $\lim_{n\rightarrow\infty}n+\frac{1}{2}-\mathscr{C}(C_{2n+1})=0$, so we conclude $\lim_{n\rightarrow\infty}4n+2-\mathscr{C}(J(4n+2,3))=0$.
3.2 Lower bound on the independence number
Given a graph $G$, the standard approach to proving lower bounds on $\alpha(G)$ is by exhibiting an independent set. For instance, in [35], Bohman computes a lower bound on the independence number of powers of an odd cycle. The structure of odd cycles is used to show that for any natural number $d$, $m\geq 3$ and even number $\beta>0$,
$\alpha(C^{d}_{m+\beta})\geq\alpha(C^{d}_{m}){\left(\dfrac{m+\beta}{m}\right)}^{d}$ (20)
where $C^{d}_{m}$ is the $d$-th power of $C_{m}$. The above inequality is non-trivial when $m$ is an odd number. Note that for any two graphs $G$ and $H$ and any natural number $d$, (7) yields $\frac{\alpha(G^{d})}{\alpha(H^{d})}\leq\alpha^{*}(G|H)^{d}.$ In Appendix C, we prove that $\alpha^{*}(C_{n}|C_{m})=n/m$ for odd $m$ and $n$ satisfying $n<m$. Thus, we recover the result of [35]. However, the more important point is that the structure of the cycle graphs is not a fundamental constraint in our calculation; we can prove a lower bound on $\alpha(H^{d})$ in terms of $\alpha(G^{d})$ if we can compute $\alpha^{*}(G|H)$ for any arbitrary graphs $G$ and $H$. In the next sub-section, we extend the result of [3] and compute $\alpha^{*}(G|H)$ for Cayley graphs.
3.3 The ratio of zero-error capacities of two graphs
By Theorem 3, the inequality $\frac{\mathscr{C}(G)}{\mathscr{C}(H)}\leq\alpha^{*}(G|H)$ yields a non-trivial upper bound on the ratio of Shannon number of $G$ and $H$. To compare this bound with the previously known bounds, observe that one can use individual bounds on $\mathscr{C}(G)$ and $\mathscr{C}(H)$ to write the following upper bound. In particular, we can write
$\frac{\mathscr{C}(G)}{\mathscr{C}(H)}\leq\frac{\mathscr{C}(G)}{(\alpha(H^{d}))^{1/d}}$ (21)
for any natural number $d$ and then use a known upper bound on $\mathscr{C}(G)$ (such as the Lovász number) to compute an upper bound on the ratio of the capacities of $G$ and $H$. However, calculating $\alpha(H^{d})$ for large values of $d$ is computationally difficult, and choosing small values of $d$ can lead to weak bounds. As an example, for a small graph like $C_{7}$, it is known that $\alpha(C^{3}_{7})=33$, and for the larger values of $d$, we only know the bounds $108\leq\alpha(C^{4}_{7})\leq 115$ and $367\leq\alpha(C^{5}_{7})\leq 401$ [32]. Also as we show in Appendix C, $\frac{\mathscr{C}(C_{9})}{\mathscr{C}(C_{11})}\leq\alpha^{*}(C_{9}|C_{11})=\frac{9}{11}$, while the best we can have using the previous capacity results is that $\mathscr{C}(C_{9})\leq 4.3601$ and $\mathscr{C}(C_{11})\geq 148^{\frac{1}{3}}>5.2895$ [33], so $\frac{\mathscr{C}(C_{9})}{\mathscr{C}(C_{11})}\leq\frac{4.3601}{5.2895}$. We have $\frac{9}{11}\leq\frac{4.3601}{5.2895}$. This illustrates the importance of directly finding upper bounds on the ratio of the Shannon capacities of graphs $G$ and $H$ rather than studying the capacities of $G$ and $H$ separately. As a concrete example, we compute $\alpha^{*}(G|H)$ for two Cayley graphs obtained by cyclic groups and obtain a novel result that (to the best of our knowledge) cannot be derived using previous results.
Theorem 9.
Let $G=Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)$ and $H=Cay(\mathbb{Z}_{m},\pm 1,\pm 2,\dots,\pm k)$, where $1\leq 2k<n<m$ are integers. Then, $\alpha^{*}(G|H)\geq\frac{n}{m}$. Moreover, $\alpha^{*}(G|H)=\frac{n}{m}$ if there are integers $\ell,s\geq 0$ such that $m=\ell n+s(k+1)$.
Proof of the above theorem is given in Appendix E. It utilizes the idea of defining a homomorphism from $H$ to $G$. Theorem 9 implies $\frac{\mathscr{C}(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k))}{\mathscr{C}(Cay(\mathbb{Z}_{m},\pm 1,\pm 2,\dots\pm k))}\leq\frac{n}{m}$ for $m=\ell n+s(k+1)$ for some integers $\ell,s\geq 0$. This inequality is an explicit and non-trivial upper bound on the ratio of the capacities of Cayley graphs. We believe this result is novel. To compare this bound with the previously known bounds, one can use individual bounds on $\mathscr{C}(G)$ and $\mathscr{C}(H)$ to write the following upper bound. In [28], it is proved that
$\alpha(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)\boxtimes Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k))=\left\lfloor\frac{\left\lfloor\frac{n}{k+1}\right\rfloor n}{k+1}\right\rfloor.$
So, $\frac{\mathscr{C}(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots k))}{\mathscr{C}(Cay(\mathbb{Z}_{m},\pm 1,\pm 2,\dots k))}\leq\frac{\alpha^{*}(Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k))}{\alpha(Cay(\mathbb{Z}_{m},\pm 1,\pm 2,\dots\pm k)^{2})}=\frac{n/(k+1)}{\sqrt{\left\lfloor\frac{\left\lfloor\frac{m}{k+1}\right\rfloor m}{k+1}\right\rfloor}}$, which is strictly greater than $\frac{n}{m}$ if $m$ is not divisible by $k+1$.
3.4 Homomorphism and the relative fractional independence number
The well-known No-Homomorphism Lemma states that [^9] See [36, Exercise 2.12] for an extension of the No-Homomorphism Lemma based on the maximum number of vertices in an induced sub-graph of $G$ that is homomorphic to an auxiliary graph $K$.:
Lemma 6.
[37] If there is a homomorphism from $H$ to $G$, and $G$ is vertex-transitive, then $\frac{\alpha(G)}{\alpha(H)}\leq\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}$.
We can use No-Homomorphism Lemma to show the non-existence of homomorphism from $H$ to $G$. Also, if we show that there is a homomorphism from $H$ to $G$, then we can use No-Homomorphism Lemma to give an upper bound on the $\alpha(G)$ (See [38]). We give an extension of the No-Homomorphism Lemma using the relative fractional independence number.
Lemma 7.
If $G$ and $H$ are two graphs where there is a homomorphism $\mathsf{g}:H\rightarrow G$ then $\alpha(G^{c}\boxtimes H)\geq|\mathcal{V}(H)|$.
Proof.
Note that the collection of vertices $(\mathsf{g}(u),u)$ in $G^{c}\boxtimes H$ for all $u\in\mathcal{V}(H)$ is an independent set of size $|\mathcal{V}(H)|$. ∎
Lemma 8.
If there is a homomorphism from $H$ to $G$, and $G$ is vertex-transitive, then $\alpha^{*}(G|H)\leq\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}$.
Proof.
Since $G$ is vertex-transitive, by Theorem 4, $\alpha^{*}(G|H)=\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)}$. By Lemma 7, since there is a homomorphism from $H$ to $G$, we have $\alpha(G^{c}\boxtimes H)\geq|\mathcal{V}(H)|$.
We have the following corollary by Lemma 8 and Theorem 3.
Corollary 7.
If there is a homomorphism from $H$ to $G$, and $G$ is vertex-transitive, then $\frac{X(G)}{X(H)}\leq\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}$, where $X(G)$ can be the zero-error Shannon capacity of $G$, the fractional independence number of $G$, the Lovász number of $G$, Schrijver’s or Szegedy’s variants of the Lovász number.
Now, we give two examples to show that our extension of the No-Homomorphism Lemma is stronger than Lemma 6.
Example 3.
Suppose that $G$ is the Petersen graph with $10$ vertices, so $\alpha(G)=4$ and $\alpha^{*}(G)=5$. Take any arbitrary graph $H$ such that $2.5\alpha(H)\geq|\mathcal{V}(H)|>2\alpha^{*}(H)$. For example, we can take $H$ as a cycle graph $C_{3}$ with a path on two vertices connected to one of the vertices of $C_{3}$. So, $H$ has $5$ vertices and $\alpha(H)=\alpha^{*}(H)=2$. By Lemma 8, there is no homomorphism from $G$ to $H$ because $\frac{\alpha^{*}(G)}{\alpha^{*}(H)}=\frac{5}{\alpha^{*}(H)}>\frac{10}{|\mathcal{V}(H)|}$. However, we cannot deduce this from the standard No-Homomorphism Lemma because $\frac{\alpha(G)}{\alpha(H)}=\frac{4}{\alpha(H)}\leq\frac{10}{|\mathcal{V}(H)|}$.
Example 4.
Suppose that $H=Cay(Z_{7},\{\pm 1,\pm 2\})$, i.e $H$ is a graph with vertex set $\{v_{1},\dots,v_{7}\}$ and $v_{i}$ is connected to $v_{j}$ if $i-j\mod 7=\pm 1$ or ±$\pm 2$, and $G=Cay(Z_{8},\{\pm 1,\pm 2\})$. Then $|\mathcal{V}(H)|=7$ and $|\mathcal{V}(G)|=8$ and $\alpha(H)=\alpha(G)=2$. There is no homomorphism from $H$ to $G$, though this can not be deduced from No-Homomorphism Lemma, because $\frac{2}{2}\leq\frac{8}{7}$. But, according to Lemma 8, if there exists a homomorphism from $H$ to $G$ then we should have $\alpha^{*}(G|H)\leq\frac{8}{7}$. We know that $\alpha^{*}(G|H)=\frac{8}{\alpha(G^{c}\boxtimes H)}$. By inspection $\alpha(G^{c}\boxtimes H)=6$ and since $8/6>8/7$, Lemma 8 implies that there exists no homomorphism from $H$ to $G$.
Finally, we also note that our results provide connections with the previously known results:
Theorem 10 (Lemma 3 in [19]).
For any vertex-transitive graph $G$ and any arbitrary graph $W$ we have
$\alpha(G\boxtimes W)=\min_{H\subset G}\left\{\alpha(H\boxtimes W)\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}\right\}$
where $H\subset G$ means that the minimum is over all graphs $H$ that are induced subgraph of $G$.
The above result shows that if $H$ is an induced subgraph of $G$ then
$\alpha^{*}(G|H)\leq\frac{|\mathcal{V}(G)|}{|\mathcal{V}(H)|}.$
We can also deduce this from our No-Homomorphism lemma (Lemma 8), because if $H$ is an induced subgraph of $G$, then there is a homomorphism from $H$ to $G$.
4 Some open problems
  • We derived the following inequalities:
    $\frac{1}{\alpha^{*}(H|G)}\leq\frac{\alpha^{*}(G|W)}{\alpha^{*}(H|W)}\leq\alpha^{*}(G|H)\qquad\forall\text{ graph }W,$
    and
    $\frac{1}{\alpha^{*}(H|G)}\leq\frac{X(G)}{X(H)}\leq\alpha^{*}(G|H)$
    where $X(\cdot)$ can be any of the properties described in Theorem 3. One might be interested in characterizing graphs $G$ and $H$ for which $\alpha^{*}(G|H)\alpha^{*}(H|G)=1$ holds, implying that all of the above inequalities hold with equality. The equality $\alpha^{*}(G|H)\alpha^{*}(H|G)=1$ holds if and only if
    $\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}=\frac{\alpha(G)}{\alpha(H)}$
    for every graph $W$. In other words, every graph $W$ should be a maximizer of $\alpha^{*}(G|H)$. An example that this may occur is as follows: suppose that
    $G=G_{1}\boxtimes G^{\prime},\qquad H=G_{1}\boxtimes H^{\prime}$
    where $G_{1}$ is some arbitrary graph, and both $G^{\prime}$ and $H^{\prime}$ are universal graphs. Then the equality $\alpha^{*}(G|H)\alpha^{*}(H|G)=1$ holds. However, we do not have a full characterization of when the equality $\alpha^{*}(G|H)\alpha^{*}(H|G)=1$ holds in general.
  • Is there any connections between $\alpha^{*}(G|H)$ and the fractional Haemers numbers $G$ and $H$?
  • Is there an efficient algorithm to check whether $G\in Expand(H)$?
  • In this paper, we considered the following expansions:
    $\alpha(G^{n})=\prod_{i=1}^{n}\frac{\alpha(G^{i})}{\alpha(G^{i-1})}=\prod_{i=1}^{n}\frac{\alpha(G\boxtimes G^{i-1})}{\alpha(G^{i-1})}\leq\left[\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}\right]^{n},$ (22)
    and
    $\frac{\alpha(G^{n})}{\alpha(H^{n})}=\prod_{i=1}^{n}\frac{\alpha(G\boxtimes G^{i-1}\boxtimes H^{n-i})}{\alpha(H\boxtimes G^{i-1}\boxtimes H^{n-i})}\leq\left[\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}\right]^{n}.$ (23)
    Next, we obtain a weak upper bound in (22) and (23) since we relax the supremum over all possible graphs $W$. This motivates the following conjecture. Proving or disproving the conjecture helps understand or improve the bounds in (22) and (23).
    Conjecture 1.
    The limit $\lim_{i\rightarrow\infty}\dfrac{\alpha(G\boxtimes G^{i-1})}{\alpha(G^{i-1})}$ exists for any graph $G$. Moreover, for any $c\in(0,1)$, the limit
    $\lim_{n\rightarrow\infty}\frac{\alpha(G\boxtimes G^{nc}\boxtimes H^{n(1-c)})}{\alpha(H\boxtimes G^{nc}\boxtimes H^{n(1-c)})}$
    exists for any graph $G$ and $H$.
5 Proofs
5.1 Proof of Theorem 1
Observe that the theorem’s second and third parts follow from the first part. For the third part, the dual of the linear program in the first part is as follows: assigning a non-negative weight $\beta_{f}$ to $f\in\mathcal{F}(H,G)$ and multiplying the equation $\sum_{i=1}^{k}w_{i}|f(v_{i})|\leq 1$ by $\beta_{f}$; we obtain the following dual to the linear program of the first part: $\alpha^{*}(G|H)$ equals the minimum of $\sum_{f}\beta_{f}$ such that
$\sum_{f}\beta_{f}|f(v)|\geq 1,\qquad\forall v\in\mathcal{V}(G).$
Equivalently, $\alpha^{*}(G|H)$ equals the minimum over $\beta_{f}$ of
$\frac{\sum_{f}\beta_{f}}{\min_{v\in\mathcal{V}(G)}\sum_{f}\beta_{f}|f(v)|}.$
By scaling $\beta_{f}$, we can impose the constraint $\sum_{f}\beta_{f}=1$ and view $\beta_{f}$ as a probability distribution over $f\in\mathcal{F}(H,G)$. Hence, the third part of the theorem follows. The third part implies the second part, as the second part implies that
$\frac{1}{\alpha^{*}(G|H)}=\max\min_{v\in\mathcal{V}(G)}\mathbb{E}[|F(v)|].$
Using the minimax theorem and exchanging the order of minimum and maximum, we have
$\frac{1}{\alpha^{*}(G|H)}=\min_{w_{v}\geq 0,\sum_{v}w_{v}=1}\max_{f}\sum_{v\in V(G)}w_{v}|f(v)|.$
It remains to prove the first and fourth parts of the theorem. Below, we provide a proof for these. An alternative proof is given in Appendix D. We begin with the following lemma.
Lemma 9.
To compute $\alpha^{*}(G|H)$, it suffices to take supremum over graphs $W$ with the following structure: the vertex set of $W$ partitions as follows: $\mathcal{V}({W})=\bigcup_{\mathcal{S}\in\mathcal{I}(G)}\mathcal{W}_{\mathcal{S}}$ for some disjoint sets $\mathcal{W}_{\mathcal{S}}$. The edge structure in $W$ is as follows: for any $\mathcal{S}_{1},\mathcal{S}_{2}\in\mathcal{I}(G)$ vertex $u_{1}\in\mathcal{W}_{S_{1}}$ is connected to $u_{2}\in\mathcal{W}_{\mathcal{S}_{2}}$ iff $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are disconnected in $G$. In particular, if $\mathcal{S}_{1}=\mathcal{S}_{2}$ and $u_{1},u_{2}\in\mathcal{W}_{\mathcal{S}}$, then $u_{1}$ and $u_{2}$ are not connected. Moreover, we have
$\alpha(W\boxtimes G)=\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|\mathcal{S}|.$
Proof.
Take some arbitrary graph $W$. Let $\mathcal{B}\subseteq\mathcal{V}(W)\times\mathcal{V}(G)$ be an independent set for $W\boxtimes G$ of maximum size. For every $u\in\mathcal{V}(W)$, let $\mathcal{B}_{u}=\{v\in\mathcal{V}(G):(u,v)\in\mathcal{B}\}.$ If $\mathcal{B}_{u}=\emptyset$ for some vertex $u$ of $W$, we can simply remove this vertex from $W$. This removal would not affect $\alpha(W\boxtimes G)$ but can potentially reduce $\alpha(W\boxtimes H)$. Therefore, without loss of generality, we can assume that $\mathcal{B}_{u}\neq\emptyset$ for all vertices $u$ of $W$.
Since $\mathcal{B}$ is an independent set for $W\boxtimes G$, for every $u\in\mathcal{V}(W)$, $\mathcal{B}_{u}$ must be an independent set for $G$. Thus, $\mathcal{B}_{u}\in\mathcal{I}(G)$. This leads to the following partition of the vertex set $\mathcal{V}(W)$: for every non-empty $\mathcal{S}\in\mathcal{I}(G)$, let $\mathcal{W}_{\mathcal{S}}=\{u\in\mathcal{V}(W):\mathcal{B}_{u}=\mathcal{S}\}.$ With this definition, the condition that $\mathcal{B}\subseteq\mathcal{V}(W)\times\mathcal{V}(H)$ is an independent set for $W\boxtimes G$ is equivalent with the following condition: take $u_{1}\in\mathcal{W}_{\mathcal{S}_{1}}$ and $u_{2}\in\mathcal{W}_{\mathcal{S}_{2}}$. If $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are not disconnected (i.e., either $\mathcal{S}_{1}\cap\mathcal{S}_{2}\neq\emptyset$ or there is an edge in $G$ between a vertex in $\mathcal{S}_{1}$ and a vertex in $\mathcal{S}_{2}$) then there should not be any edge between $u_{1},u_{2}$ in $W$. If $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are disconnected; there may or may not be edges between vertices in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. We claim that, without loss of generality, we can add edges between the vertices in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ if $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are disconnected since their addition would not affect $\alpha(W\boxtimes G)$ but could potentially decrease $\alpha(W\boxtimes H)$, which is desirable as we want to maximize the ratio $\alpha(W\boxtimes G)/\alpha(W\boxtimes H)$. Thus, the graph $W$ will have the form given in the lemma statement. We also have
$|\mathcal{B}|=\sum_{u\in\mathcal{V}(W)}|\mathcal{B}_{u}|=\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|\mathcal{S}|.$
Consider a graph $W$ with the structure in Lemma 9. We claim that
$\alpha(W\boxtimes H)=\max_{f\in\mathcal{F}}\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|f(\mathcal{S})|$ (24)
where $\mathcal{F}$ is the set of all functions $f:\mathcal{I}(G)\mapsto\mathcal{I}(H)$ such that for any distinct $\mathcal{S}_{1},\mathcal{\mathcal{missing}}S_{2}\in\mathcal{I}(G)$, if $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are disconnected in $G$ then $f(\mathcal{S}_{1})$ and $f(\mathcal{S}_{2})$ are disconnected in $H$. First, observe that $\alpha(W\boxtimes H)\geq\max_{f\in\mathcal{F}}\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|f(\mathcal{S})|.$ To see this, for every function $f\in\mathcal{F}$, $\{(u,v)\big{|}u\in\mathcal{W}_{\mathcal{S}},v\in f(\mathcal{S})\text{ for some}\mathcal{S}\in\mathcal{I}(G)\}$ is an independent set. On the other hand, take an independent set $\mathcal{A}$ for $W\boxtimes H$ of maximum size. Take some $\mathcal{S}\in\mathcal{I}(G)$ and for every $u\in\mathcal{\mathcal{W}}_{\mathcal{S}}$ consider $\mathcal{A}_{u}=\{v\in\mathcal{V}(H):(u,v)\in\mathcal{A}\}.$ Since $\mathcal{A}$ is an independent set for $W\boxtimes H$, we have $\mathcal{A}_{u}\in\mathcal{I}(H)$. Consider $|\mathcal{A}_{u}|$ for $u\in\mathcal{W}_{\mathcal{S}}$, and let $u^{*}\in\mathcal{W}_{\mathcal{S}}$ have maximum $|\mathcal{A}_{u^{*}}|$. For all $u\in\mathcal{W}_{\mathcal{S}}$, let us replace $\mathcal{A}_{u}$ by $\mathcal{A}_{u^{*}}$. This transformation will not decrease the size of $\mathcal{A}$ (as $|\mathcal{A}_{u^{*}}|$ had maximum size) and will keep $\mathcal{A}$ as an independent set for $W\boxtimes H$. If we make this transformation on all $\mathcal{S}\in\mathcal{I}(G)$, we can force all $u\in\mathcal{W}_{\mathcal{S}}$ mapped to the same independent set in $\mathcal{I}(H)$, which we can call $f(\mathcal{S})\in\mathcal{I}(H)$. Thus, the size of the independent set $\mathcal{A}$ is no more than $\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|f(\mathcal{S})|$. Hence, $\alpha(W\boxtimes H)\leq\max_{f\in\mathcal{F}}\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|f(\mathcal{S})|.$ This implies (24). To sum this up, we obtain
$\frac{\alpha(W\boxtimes G)}{\alpha(W\boxtimes H)}=\frac{\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|\mathcal{S}|}{\max_{f\in\mathcal{F}}\sum_{\mathcal{S}\in\mathcal{I}(G)}|\mathcal{W}_{\mathcal{S}}|\times|f(\mathcal{S})|}.$
Observe that $|\mathcal{W}_{\mathcal{S}}|$ can be set to any arbitrary natural number. Letting $\omega_{\mathcal{S}}=|\mathcal{W}_{\mathcal{S}}|$ we obtain
$\sup_{W}\frac{\alpha(W\boxtimes G)}{\alpha(W\boxtimes H)}=\sup_{\{\omega_{\mathcal{S}}\in\mathbb{Z}\cup\{0\}:\mathcal{S}\in\mathcal{I}(G)\}}~{}\min_{f\in\mathcal{F}}\frac{\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega_{\mathcal{S}}|\mathcal{S}|}{\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega_{\mathcal{S}}|f(S)|}.$
In particular, if the supremum on the right-hand side is a maximum, the supremum on the left-hand side is also a maximum. Next, we claim that to compute
$\sup_{\{\omega_{\mathcal{S}}\in\mathbb{Z}\cup\{0\}:\mathcal{S}\in\mathcal{I}(H)\}}~{}\min_{f\in\mathcal{F}}\frac{\sum_{\mathcal{S}\in\mathcal{I}(H)}\omega_{\mathcal{S}}|\mathcal{S}|}{\sum_{\mathcal{S}\in\mathcal{I}(H)}\omega_{\mathcal{S}}|f(\mathcal{S})|},$ (25)
without loss of generality, one can assume that $\omega_{\mathcal{S}}=0$ if $|\mathcal{S}|\neq 1$. We first show that without loss of generality, to compute (25), we can restrict the minimization over $f\in\mathcal{F}$ satisfying
$|f(\mathcal{S})|\geq\sum_{i\in\mathcal{S}}|f(\{i\})|,$
for any set $\mathcal{S}\in\mathcal{I}(G)$. To show this, take an arbitrary function $f$ and $\mathcal{S}\in\mathcal{I}(G)$ such that
$|f(\mathcal{S})|<\sum_{i\in\mathcal{S}}|f(\{i\})|.$
Observe that $f(\{i\})$ and $f(\{j\})$ for $i,j\in\mathcal{S}$ are disconnected independent sets in $\mathcal{I}(H)$, since $\{i\}$ and $\{j\}$ are disconnected independent sets in $\mathcal{I}(G)$. Thus,
$\sum_{i\in\mathcal{S}}|f(\{i\})|=\big{|}\cup_{i\in\mathcal{S}}f(\{i\})\big{|}.$
We claim that replacing $f(\mathcal{S})$ by $\cup_{i\in\mathcal{S}}f(\{i\})$ would not violate conditions on $f$. This change would increase $|f(\mathcal{S})|$ and hence would also increase the term $\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega_{\mathcal{S}}\cdot|f(\mathcal{S})|$. To show that the replacement would not violate conditions on $f$, take some arbitrary set $\mathcal{T}\in\mathcal{I}(G)$ where $\mathcal{T}$ and $\mathcal{S}$ are disconnected in $G$. Then, the set $\{i\}$ and $\mathcal{T}$ are disconnected in $G$, implying that $f(\mathcal{T})\in\mathcal{I}(H)$ is disconnected from $f(\{i\})$. Hence, $f(\mathcal{T})$ is disconnected from $\cup_{i\in\mathcal{S}}f(\{i\})$. Thus, the desired condition is satisfied by this replacement. To sum this up, without loss of generality, we can restrict the minimization over $f$ to those satisfying
$|f(\mathcal{S})|\geq\sum_{i\in\mathcal{S}}|f(\{i\})|.$
Now, take some arbitrary set of $\{\omega_{\mathcal{S}}\geq 0:\mathcal{S}\in\mathcal{I}(G)\}$. Then, consider the following:
$\omega^{\prime}(\{i\})$ $=\sum_{\mathcal{S}:i\in\mathcal{S}}\omega_{\mathcal{S}}.$ (26)
$\omega^{\prime}(\mathcal{S})$ $=0\qquad\text{ if }|\mathcal{S}|>1.$ (27)
Then, we have
$\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega^{\prime}_{\mathcal{S}}\cdot|S|$ $=\sum_{i}\sum_{\mathcal{S}:~{}i\in\mathcal{S}}\omega_{\mathcal{S}}$ (28)
$=\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega_{\mathcal{S}}\cdot|\mathcal{S}|$ (29)
and
$\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega^{\prime}_{\mathcal{S}}\cdot|f(\mathcal{S})|$ $=\sum_{i}\sum_{\mathcal{S}:~{}i\in\mathcal{S}}\omega_{\mathcal{S}}\cdot|f(\{i\})|$ (30)
$=\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega_{\mathcal{S}}\sum_{i\in\mathcal{\mathcal{missing}}S}|f(\{i\})|$ (31)
$\leq\sum_{\mathcal{S}\in\mathcal{I}(G)}\omega_{S}\cdot|f(\mathcal{S})|.$ (32)
Since this holds for every arbitrary function $f$, this implies that without loss of generality, we can assume that $\omega_{\mathcal{S}}$ is non-zero only when $|\mathcal{S}|=1$. Assuming that $\mathcal{V}(G)=\{v_{1},v_{2},\cdots,v_{k}\}$, let $w_{i}=\omega_{\{i\}}$, and $\mathcal{T}_{i}=f(\{i\})$. Then,
$\sup_{W}\frac{\alpha(W\boxtimes G)}{\alpha(W\boxtimes H)}=\sup_{\{w_{i}\in\mathbb{Z}\cup\{0\}\}}~{}\min_{\mathcal{T}_{1},\cdots,\mathcal{T}_{k}}\frac{\sum_{i=1}^{k}w_{i}}{\sum_{i=1}^{k}w_{i}|\mathcal{T}_{i}|}$
where the minimum is over any collection of sets $\mathcal{T}_{1},\cdots,\mathcal{T}_{k}\in\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected in $H$ if there is no edge between $v_{i}$ and $v_{j}$ in $G$.
Observe that scaling $w_{i}$ would not change the above expression. Therefore,
$\sup_{W}\frac{\alpha(W\boxtimes G)}{\alpha(W\boxtimes H)}=\sup_{\{w_{i}\in\mathbb{Q},w_{i}\geq 0\}}~{}\sum_{i=1}^{k}w_{i}$
subject to $\sum_{i=1}^{k}w_{i}|\mathcal{T}_{i}|\leq 1$ for any collection of sets $\mathcal{T}_{1},\cdots,\mathcal{T}_{k}\in\mathcal{I}(H)$ where $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected if $v_{i}v_{j}\notin\mathcal{E}(G)$. Relaxing $w_{i}\in\mathbb{Q}$ to $w_{i}\in\mathbb{R}$ would not change the value of this linear program because the solution of this linear program is rational. The first part of the theorem thus follows. To show the last part, observe that the domain of the linear program is compact since $w_{i}\leq 1$ for all $i$. To see this, take $\mathcal{T}_{i}$ as an independent set in $H$ and $\mathcal{T}_{j}=\emptyset$ for $j\neq i$. The compactness of the domain of the linear program implies that the solution is obtained at some finite $(w_{1},\cdots,w_{k})$. Thus, the supremum in the definition of $\alpha^{*}(G|H)$ is a maximum.
5.2 Proof of Theorem 2
For the first part of the theorem, observe that
$\alpha^{*}(G_{1}\boxtimes G_{2}|H_{1}\boxtimes H_{2})$ $=\sup_{W}\frac{\alpha(W\boxtimes G_{1}\boxtimes G_{2})}{\alpha(W\boxtimes H_{1}\boxtimes H_{2})}=\sup_{W}\frac{\alpha(W\boxtimes G_{1}\boxtimes G_{2})}{\alpha(W\boxtimes H_{1}\boxtimes G_{2})}\cdot\frac{\alpha(W\boxtimes H_{1}\boxtimes G_{2})}{\alpha(W\boxtimes H_{1}\boxtimes H_{2})}$
$\leq\sup_{W}\frac{\alpha((W\boxtimes G_{2})\boxtimes G_{1})}{\alpha((W\boxtimes G_{2})\boxtimes H_{1})}\cdot\sup_{W}\frac{\alpha((W\boxtimes H_{1})\boxtimes G_{2})}{\alpha((W\boxtimes H_{1})\boxtimes H_{2})}$
$\leq\alpha^{*}(G_{1}|H_{1})\alpha^{*}(G_{2}|H_{2}).$
For the second part, let $C$ be the maximizer for $\alpha^{*}(G|W)$. Then, we have
$\frac{\alpha^{*}(G|W)}{\alpha^{*}(H|W)}=\frac{\sup_{A}\frac{\alpha(G\boxtimes A)}{\alpha(W\boxtimes A)}}{\sup_{B}\frac{\alpha(H\boxtimes B)}{\alpha(W\boxtimes B)}}=\frac{\frac{\alpha(G\boxtimes C)}{\alpha(W\boxtimes C)}}{\sup_{B}\frac{\alpha(H\boxtimes B)}{\alpha(W\boxtimes B)}}\leq\frac{\frac{\alpha(G\boxtimes C)}{\alpha(W\boxtimes C)}}{\frac{\alpha(H\boxtimes C)}{\alpha(W\boxtimes C)}}=\frac{\alpha(G\boxtimes C)}{\alpha(H\boxtimes C)}\leq\alpha^{*}(G|H).$
Similarly, let $D$ be the maximizer for $\alpha^{*}(H|W)$. We have
$\frac{\alpha^{*}(G|W)}{\alpha^{*}(H|W)}=\frac{\sup_{A}\frac{\alpha(G\boxtimes A)}{\alpha(W\boxtimes A)}}{\sup_{B}\frac{\alpha(H\boxtimes B)}{\alpha(W\boxtimes B)}}=\frac{\sup_{A}\frac{\alpha(G\boxtimes A)}{\alpha(W\boxtimes A)}}{\frac{\alpha(H\boxtimes D)}{\alpha(W\boxtimes D)}}\geq\frac{\frac{\alpha(G\boxtimes D)}{\alpha(W\boxtimes D)}}{\frac{\alpha(H\boxtimes D)}{\alpha(W\boxtimes D)}}=\frac{\alpha(G\boxtimes D)}{\alpha(H\boxtimes D)}\geq\frac{1}{\alpha^{*}(H|G)}.$
For the third part, consider the max form of the linear programs (first part of Theorem 1) for computing $\alpha^{*}(G_{1}+G_{2}|H)$, $\alpha^{*}(G_{1}|H)$ and $\alpha^{*}(G_{2}|H)$ and let us denote the set of constraints in these linear programs by $Cons(G_{1}+G_{2}|H),Cons(G_{1}|H)$ and $Cons(G_{2}|H)$, respectively. Thus, for instance, $\alpha^{*}(G_{1}+G_{2}|H)=\max\sum_{v_{i}\in\mathcal{V}(G_{1})}w_{i}+\sum_{v_{i}\in\mathcal{V}(G_{2})}w_{i}$ subject to $Cons(G_{1}+G_{2})$. Also $\alpha^{*}(G_{1}|H)=\max\sum_{v_{i}\in\mathcal{V}(G_{1})}w_{i}$ subject to $Cons(G_{1}|H)$ and $\sum_{v_{i}\in\mathcal{V}(G_{2})}w_{i}$ subject to $Cons(G_{2}|H)$. It is clear that, $Cons(G_{1}|H)\subset Cons(G_{1}+G_{2}|H)$ and $Cons(G_{2}|H)\subset Cons(G_{1}+G_{2}|H)$, so $\alpha^{*}(G_{1}+G_{2}|H)\leq\alpha^{*}(G_{1}|H)+\alpha^{*}(G_{2}|H).$
For the forth part, it is clear that $\alpha^{*}((G_{1}+G_{2})^{c}|H)\geq\max(\alpha^{*}(G_{1}^{c}|H),\alpha^{*}(G_{2}^{c}|H))$. Consider the characterization in the third part of Theorem 1 to show the reverse direction. Take the optimal random mappings for $\alpha^{*}(G_{1}^{c}|H)$ and $\alpha^{*}(G_{2}^{c}|H)$ and combine them to construct a random mapping for $(G_{1}+G_{2})^{c}$. This random mapping implies that
$\alpha^{*}((G_{1}+G_{2})^{c}|H)\leq\max(\alpha^{*}(G_{1}^{c}|H),\alpha^{*}(G_{2}^{c}|H)).$
For the fifth part, we have
$\frac{1}{\alpha^{*}(G|H_{1}+H_{2})}$ $=\inf_{W}\frac{\alpha((H_{1}+H_{2})\boxtimes W)}{\alpha(G\boxtimes W)}=\inf_{W}\frac{\alpha(H_{1}\boxtimes W)+\alpha(H_{2}\boxtimes W)}{\alpha(G\boxtimes W)}$
$\geq\inf_{W}\frac{\alpha(H_{1}\boxtimes W)}{\alpha(G\boxtimes W)}+\inf_{W}\frac{\alpha(H_{2}\boxtimes W)}{\alpha(G\boxtimes W)}=\frac{1}{\alpha^{*}(G|H_{1})}+\frac{1}{\alpha^{*}(G|H_{2})}.$
According to Theorem 4, when $G$ is a vertex-transitive graph on $k$ vertices, we have the following:
$\inf_{W}\frac{\alpha((H_{1}+H_{2})\boxtimes W)}{\alpha(G\boxtimes W)}$ $=\frac{\alpha((H_{1}+H_{2})\boxtimes G^{c})}{k}=\frac{\alpha(H_{1}\boxtimes G^{c})}{k}+\frac{\alpha(H_{2}\boxtimes G^{c})}{k}$
$=\inf_{W}\frac{\alpha(H_{1}\boxtimes W)}{\alpha(G\boxtimes W)}+\inf_{W}\frac{\alpha(H_{2}\boxtimes W)}{\alpha(G\boxtimes W)}=\frac{1}{\alpha^{*}(G|H_{1})}+\frac{1}{\alpha^{*}(G|H_{2})}.$
Thus, equality holds in this case.
5.3 Proof of Lemma 1
Take some $W$ such that $\alpha^{*}(G_{1}\boxtimes G_{2}\boxtimes\cdots\boxtimes G_{r})=\frac{\alpha(G_{1}\boxtimes G_{2}\boxtimes\cdots\boxtimes G_{r}\boxtimes W)}{\alpha(W)}$.
We show that this choice of $W$ works for us. Note that
$\prod_{i=1}^{r}\alpha^{*}(G_{i})$ $=\alpha^{*}(G_{1}\boxtimes G_{2}\boxtimes\cdots\boxtimes G_{r})=\frac{\alpha(G_{1}\boxtimes G_{2}\boxtimes\cdots\boxtimes G_{r}\boxtimes W)}{\alpha(W)}$
$=\frac{\alpha(G_{2}\boxtimes\cdots\boxtimes G_{r}\boxtimes G_{1}\boxtimes W)}{\alpha(G_{1}\boxtimes W)}\cdot\frac{\alpha(G_{1}\boxtimes W)}{\alpha(W)}$
$\leq\max_{W^{\prime}}\frac{\alpha(G_{2}\boxtimes\cdots\boxtimes G_{r}\boxtimes W^{\prime})}{\alpha(W^{\prime})}\cdot\max_{W^{\prime}}\frac{\alpha(G_{1}\boxtimes W^{\prime})}{\alpha(W^{\prime})}$ (33)
$=\alpha^{*}(G_{2}\boxtimes\cdots\boxtimes G_{r})\alpha^{*}(G_{1})=\prod_{i=1}^{r}\alpha^{*}(G_{i}).$
Therefore, equality must hold in (33). Hence $\alpha^{*}(G_{1})=\frac{\alpha(G_{1}\boxtimes W)}{\alpha(W)}.$ A similar argument establishes the desired equality for $G_{i}$, $i>1$.
5.4 Proof of Theorem 3
It suffices to show that
$\alpha^{*}(G|H)\geq\frac{X(G)}{X(H)}.$ (34)
The other inequality $\frac{1}{\alpha^{*}(H|G)}\leq\frac{X(G)}{X(H)}$ is equivalent with $\alpha^{*}(H|G)\geq\frac{X(H)}{X(G)}$, which is the same inequality as in (34) with $G$ and $H$ swapped.
Observe that by choosing $W$ as a trivial graph with just one node and no edges, we have
$\alpha^{*}(G|H)=\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}\geq\frac{\alpha(G)}{\alpha(H)}.$
Next, assume that we have two quantities $X(\cdot)$ and $Y(\cdot)$ on a graph such that
$\alpha(G\boxtimes W)$ $\leq X(G)Y(W)$ (35)
for any two graphs $G$ and $W$. Then, we have
$\alpha^{*}(G|H)$ $=\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}\geq\sup_{W}\frac{\alpha(G\boxtimes W)}{X(H)Y(W)}=\frac{1}{X(H)}\sup_{W}\frac{\alpha(G\boxtimes W)}{Y(W)}.$
Consider the choice of $X(G)=Y(G)=\vartheta(G)$, the Lovász number of $G$. Then (35) is satisfied. Moreover, as shown in [39, Theorem 2],
$\sup_{W}\frac{\alpha(G\boxtimes W)}{\vartheta(W)}=\vartheta(G).$
Thus, the desired inequality holds for the Lovász number of $G$.
Similarly, the choices of $X(G)=\vartheta^{-}(G),Y(G)=\vartheta^{+}(G)$ and $X(G)=\vartheta^{+}(G),Y(G)=\vartheta^{-}(G)$. Then, by [39, Lemma 7], (35) is satisfied. From [39, Theorem 8], we have
$\sup_{W}\frac{\alpha(G\boxtimes W)}{\vartheta^{-}(W)}=\vartheta^{+}(G),$
$\sup_{W}\frac{\alpha(G\boxtimes W)}{\vartheta^{+}(W)}=\vartheta^{-}(G).$
So, we obtain the desired results for the Schrijver’s or Szegedy’s variants of the Lovász number.
5.5 Proof of Theorem 7
Note that $\mathcal{S}$ is an independent set of $(G^{\prime})^{c}\boxtimes H^{\prime}$. For any vertex $u$ in $G^{\prime}$, define
$\mathcal{A}_{u}:=\{v\in\mathcal{V}(H^{\prime}):(u,v)\in\mathcal{S}\}$
We similarly define $\mathcal{B}_{v}$ for vertices $v$ in $H^{\prime}$ as
$\mathcal{B}_{v}:=\{u\in\mathcal{V}(G^{\prime}):(u,v)\in\mathcal{S}\}.$
By assumption about $G^{\prime}$ and $H^{\prime}$, the sets $A_{u}$ and $B_{v}$ are non-empty for any $u\in G^{\prime}$ and $v\in H^{\prime}$. For any vertex $v$ of $H^{\prime}$, $\mathcal{B}_{v}$ is an independent set in $(G^{\prime})^{c}$. Therefore, it forms a clique in $G^{\prime}$. To construct $G^{\prime}$ from $H^{\prime}$ by applying the expansion operations, we first replace every vertex $v$ in $H^{\prime}$ with a clique of size $|\mathcal{B}_{v}|$ to obtain $H^{\prime}_{1}$. We can imagine the vertices of these cliques labeled by the members of $\mathcal{B}_{v}$.
If two vertices in $H^{\prime}_{1}$ labeled by $u_{i}$ and $u_{j}$ are connected by an edge in $H^{\prime}_{1}$, then either for a vertex $v$ in $H^{\prime}$ we have $u_{i},u_{j}\in\mathcal{B}_{v}$, or there exist two connected vertices $v_{i},v_{j}\in H^{\prime}$ such that $u_{i}\in\mathcal{B}_{v_{i}}$ and $u_{j}\in\mathcal{B}_{v_{j}}$. Since $\mathcal{B}_{v}$ for $v\in\mathcal{V}(H^{\prime})$ is obtained from an independent set of $(G^{\prime})^{c}\boxtimes H^{\prime}$, it follows that in both cases, $u_{i}$ must be disconnected from $u_{j}$ in $(G^{\prime})^{c}$. Consequently, there is an edge between them in $G^{\prime}$. To sum this up, if there is an edge between two vertices labeled by $u_{i}$ and $u_{j}$ in $H^{\prime}_{1}$, then there is an edge between $u_{i}$ and $u_{j}$ in $G^{\prime}$.
Finally, using Remark 5, we can merge all the vertices in $H^{\prime}_{1}$ that have the same label to obtain $H^{\prime\prime}$. The assumption that $\mathcal{A}_{u}$ is non-empty would imply that the vertices of $H^{\prime\prime}$ cover all vertices of $G^{\prime}$. This yields a graph $H^{\prime\prime}$ which is a spanning subgraph of $G^{\prime}$. Now, by adding the missing edges of $G^{\prime}$ to $H^{\prime\prime}$, we can finish the construction of $G^{\prime}$.
5.6 Proof of Theorem 4
We establish the following statement: for any graphs $G$ and $H$, where $\mathcal{V}(G)=\{v_{1},\dots,v_{k}\}$, we have
$\alpha^{*}(G|H)\geq k\min\frac{1}{\sum_{i=1}^{k}|\mathcal{T}_{i}|}=\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)},$ (36)
where the minimum is over all collection of sets $\mathcal{T}_{1},\cdots,\mathcal{T}_{k}\in\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected in $H$ if there is no edge between $v_{i}$ and $v_{j}$ in $G$. Moreover, the lower bound in (36) is tight (holds with equality) if $G$ is vertex-transitive.
According to Theorem 1, $\alpha^{*}(G|H)=\max_{\mathbf{w}}\sum_{i=1}^{k}w_{i}$, where the maximum is over all non-negative weights $w_{i}$ such that $\sum w_{i}|\mathcal{T}_{i}|\leq 1$ for all collection $\mathcal{T}_{1},\dots,\mathcal{T}_{k}\in\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected if $v_{i}v_{j}\not\in\mathcal{E}(G)$. If we only restrict to those weights that $w_{1}=\dots=w_{k}$ we get
$\alpha^{*}(G|H)\geq k\max w_{1}$
where $w_{1}\leq 1/(\sum_{i}|\mathcal{T}_{i}|)$. Since $\max w_{1}=\min 1/(\sum_{i}|\mathcal{T}_{i}|)$ the lower bound is established. Next, we show that $\max\sum_{i=1}^{k}|\mathcal{T}_{i}|=\alpha(G^{c}\boxtimes H)$. Let $\mathcal{T}^{*}_{i}$ be a valid assignment of independent sets of $H$ to the vertices of $G$ such that $\sum_{i=1}^{k}|\mathcal{T}^{*}_{i}|=\max\sum_{i=1}^{k}|\mathcal{T}_{i}|$. Consider all the pairs of vertices $v_{i},u_{j}$, such that $v_{i}\in\mathcal{V}(G)$ and $u_{j}\in\mathcal{T}^{*}_{i}$. The set of these pairs are independent in $G^{c}\boxtimes H$, so, $\max\sum_{i=1}^{k}|\mathcal{T}_{i}|\leq\alpha(G^{c}\boxtimes H)$. On the other hand, if $(v^{*}_{i},u^{*}_{j})$’s are the set of vertices in a maximum independent set, $\mathcal{I}^{*}$, of $G^{c}\boxtimes H$, then if we define $\mathcal{T}_{i}\triangleq\{u^{*}_{j}|(v^{*}_{i},u^{*}_{j})\in\mathcal{I}^{*}\}$, we get a valid assignment. This shows $\max\sum_{i=1}^{k}|\mathcal{T}_{i}|\geq\alpha(G^{c}\boxtimes H)$. Thus, $\max\sum_{i=1}^{k}|\mathcal{T}_{i}|=\alpha(G^{c}\boxtimes H)$.
Next, assume that $G$ is vertex-transitive. For any pair of vertices $v_{i}$ and $v_{j}$, let $\Pi_{i,j}$ be the set of automorphisms $\phi:G\rightarrow G$ that maps vertex $v_{i}$ to vertex $v_{j}$. Since $G$ is vertex-transitive, $|\Pi_{i,j}|\geq 1$. We claim that $|\Pi_{i,j}|=|\Pi_{1,1}|$ does not depend on $i,j$. To see this, take some $\phi^{*}_{1}\in\Pi_{1,i}$ and $\phi^{*}_{2}\in\Pi_{j,1}$. Then, $\phi\in\Pi_{i,j}$ if and only if $\phi^{*}_{2}\circ\phi\circ\phi^{*}_{1}\in\Pi_{1,1}$. This provides a one-to-one map between $\Pi_{i,j}$ and $\Pi_{1,1}$ and the claim follows.
Let $\pi$ be a permutation on $\{1,\dots,k\}$ such that $\phi(v_{i})=v_{\pi(i)}$ is an automorphism of $G$. If $\mathcal{T}_{1},\dots,\mathcal{T}_{k}$ is a collection of independent subsets of $\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected if $v_{i}v_{j}\not\in\mathcal{E}(G)$, then the constraint $\sum w_{\pi(i)}|\mathcal{T}_{i}|\leq 1$ must hold as well. Now if we take the average of all the inequalities $\sum w_{\pi(i)}|\mathcal{T}_{i}|\leq 1$ over all automorphisms (over all $\pi$), we get $\bar{w}(\sum_{i}|\mathcal{T}_{i}|)\leq 1$, where $\bar{w}=(w_{1}+\dots+w_{k})/k$. So
$\alpha^{*}(G|H)=\max_{w}(w_{1}+\dots+w_{k})=k\max\bar{w}\leq k\frac{1}{\sum_{i}|\mathcal{T}_{i}|}$
for any valid collection $\mathcal{T}_{1},\dots,\mathcal{T}_{k}\in\mathcal{I}(H)$, so $\alpha^{*}(G|H)\leq k\min\frac{1}{\sum_{i}|\mathcal{T}_{i}|}$ and the lemma holds.
5.7 Proof of Theorem 8
Assume that $\alpha^{*}(G|H)\leq 1$. We aim to prove that $G\in Expand(H)$.
5.7.1 Proof of the first part
Since $G$ is a Cayley graph, it is vertex-transitive. Therefore, by Theorem 4, we have:
$\alpha^{*}(G|H)=\frac{|\mathcal{V}(G)|}{\alpha(G^{c}\boxtimes H)}$
Thus, $\alpha^{*}(G|H)\leq 1$ implies that $|\mathcal{V}(G)|\leq\alpha(G^{c}\boxtimes H)$. Let $\mathcal{S}$ be a maximum independent set $G^{c}\boxtimes H$. For any vertex $u$ in $G$, define
$\mathcal{A}_{u}:=\{v\in\mathcal{V}(H):(u,v)\in\mathcal{S}\}$
We similarly define $\mathcal{B}_{v}$ for vertices $v$ in $H$ as
$\mathcal{B}_{v}:=\{u\in\mathcal{V}(G):(u,v)\in\mathcal{S}\}.$
We have
$\alpha(G^{c}\boxtimes H)=\sum\limits_{u\in G}|\mathcal{A}_{u}|\geq|\mathcal{V}(G)|.$
If $|A_{u}|\geq 1$ for all $u\in G$, we can apply Theorem 7 to deduce that we can construct $G$ from $H$ by applying the expansion operations. Thus, assume that $|A_{u}|=0$ for some $u\in G$. Label vertices of $G=Cay(\mathbb{Z}_{n},\pm 1,\pm 2,\dots\pm k)$ by $u_{1},u_{2},\cdots,u_{n}$ where either $n=|\mathcal{V}(G)|=c(k+1)+1$ or $n=|\mathcal{V}(G)|=c(k+1)$ for some integer $c$. Without loss of generality assume $|A_{u_{n}}|=0$. Now, consider the following $k+1$ sets in $G$:
$\mathcal{S}_{i}=\{u_{t}\,|\,t\equiv i\pmod{k+1}),\,\,t\neq n\},\qquad 1\leq i\leq k+1.$
We have $\bigcup_{i=1}^{k+1}\mathcal{S}_{i}=\mathcal{V}(G)-u_{n}$. Each set $\mathcal{S}_{i}$ is an independent set in $G$.
Since $\sum\limits_{u\in G}|\mathcal{A}_{u}|\geq n$, we must have $\sum\limits_{u\in\mathcal{S}_{i^{*}}}|\mathcal{A}_{u}|\geq\ \lceil\frac{n}{k+1}\rceil$ for some integer $11\leq i^{*}\leq k+$ 1. We construct $G$ from $H$ by applying the expansion operations as follows: we first delete every vertex $v$ from $H$ such that $\mathcal{B}_{v}\cap S_{i^{*}}=\emptyset$ to obtain $H^{\prime}$. Since $S_{i^{*}}$ was an independent set in $G$, it forms a clique in $G^{c}$. Since the projections $\mathcal{B}_{v}$ are coming from an independent set of $G^{c}\boxtimes H$, all the remaining vertices in $H^{\prime}$ must be disconnected from each other, and $H^{\prime}$ should not have any edges.
Next, due to our assumption about the set $S_{i^{*}}$, $H^{\prime}$ has at least $\lceil\frac{n}{k+1}\rceil$ vertices. To obtain $H^{\prime\prime}$, we replace $\lfloor\frac{n}{k+1}\rfloor$ of these vertices with cliques of size $k+1$. Thus, if $n=c(k+1)$, the graph $H^{\prime\prime}$ will consist of $c=\lfloor\frac{n}{k+1}\rfloor$ disjoint cliques of size $k+1$. If $n=c(k+1)+1$, the graph $H^{\prime\prime}$ will consist of $c$ disjoint cliques of size $k+1$, plus one single vertex. Hence, $H^{\prime\prime}$ is a spanning subgraph of $G$ in either case. By adding the missing edges of $H^{\prime\prime}$ to it, we complete the construction of $G$. Thus, the theorem is proved for this case.
5.7.2 Proof of the second part
Next, we show that the conjecture holds for perfect graphs. Assume that $G$ is a perfect graph with $n$ vertices and $\alpha^{*}(G|H)\leq 1$. We aim to prove that $G\in Expand(H)$. Since from Theorem 4 we have
$\alpha^{*}(G|H)\geq\frac{|\mathcal{V}(G)|}{\alpha(H\boxtimes G^{c})}$
we obtain $|\mathcal{V}(G)|\leq\alpha(H\boxtimes G^{c})$. By the Perfect Graph Theorem, we know that $G^{c}$ is also perfect. Therefore, we have $|\mathcal{V}(G)|\leq\alpha(H\boxtimes G^{c})=\alpha(H)\alpha(G^{c})$. Let $\mathcal{S}$ be a maximum independent set of $H\boxtimes G^{c}$ such that $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are maximum independent sets of $H$ and $G^{c}$ respectively.
First, we delete every vertex $u\notin\mathcal{S}_{1}$ from $H$. Secondly, replace the remaining vertices of $H$ with a clique of size $\alpha(G^{c})$ to obtain $H^{\prime}$. Due to the selection of a maximum independent set, $H^{\prime}$ consists of $\alpha(H)$ disjoint cliques, each of size $\alpha(G^{c})$. Since $\alpha^{*}(G|H)\leq 1$, we should have ${\alpha(G)}/{\alpha(H)}\leq 1$, therefore $\alpha(G)\leq\alpha(H)$ and $H^{\prime}$ has at least that $\alpha(G)$ disjoint cliques.
In Lemma 10 (provided below), we show that the vertices of a perfect graph $G$ can be partitioned into $\alpha(G)$ disjoint sets such that the induced subgraph on each set is a clique. Since cliques of $G$ are independent sets of $G^{c}$, the cliques have a maximum size of $\alpha(G^{c})$. Noting that $H^{\prime}$ consists of at least $\alpha(G)$ disjoint cliques, we can remove some vertices from $H^{\prime}$ to obtain $H^{\prime\prime}$ such that $H^{\prime\prime}$ precisely represents the clique partition of $G$. This implies that $H^{\prime\prime}$ is a spanning subgraph of $G$, and by adding the remaining edges to $H^{\prime\prime}$, we can complete the construction of $G$ from $H$.
Lemma 10.
The vertices of a perfect graph $G$ can be partitioned into $\alpha(G)$ disjoint sets such that the induced subgraph on each set is a clique.
Proof.
Since $G$ is perfect, by the Perfect Graph Theorem, $G^{c}$ is also perfect. This means the maximum clique size of $G^{c}$, which is the same as $\alpha(G)$, is equal to its chromatic number. A coloring of $G^{c}$ is equivalent to partitioning the vertices of $G^{c}$ into independent sets, and equivalently vertices of $G$ into cliques. Because $\chi(G^{c})=\alpha(G)$, this partition consists of $\alpha(G)$ cliques. Hence, the lemma is proven. ∎
6 Acknowledgment
The authors want to thank Dr. Omid Etesami for constructive discussions.
Appendix A
For example, we compute the value of $\alpha^{*}(C_{7}|C_{7}^{c})$. Let $v_{1},\dots,v_{7}$ be the vertices of $C_{7}$ and $u_{1},\dots u_{7}$ be the vertices of $C_{7}^{c}$. So $\mathcal{I}(C_{7}^{c})$ has 15 elements:
[…]
$\{u_{1},u_{2}\},\{u_{2},u_{3}\},\{u_{3},u_{4}\},\{u_{4},u_{5}\},\{u_{5},u_{6}\},\{u_{6},u_{7}\},\{u_{7},u_{1}\},\emptyset\bigg{\}}.$
Therefore, the linear program in the first part of the theorem becomes:
$\alpha^{*}(C_{7}|C_{7}^{c})=\max_{\mathbf{w}}\sum_{i=1}^{7}w_{i},$
where the maximization over $\mathbf{w}=(w_{1},w_{2},\cdots,w_{7})$ subject to $w_{i}\geq 0$ and some linear constraints on $w_{i}$’s. To write these linear constraints, we should consider all assignments of sets $\mathcal{T}_{i}\in\mathcal{I}(C_{7}^{c})$ to vertex $v_{i}\in\mathcal{V}(G)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected in $H$ if there is no edge between $v_{i}$ and $v_{j}$ in $G$.
Let us enumerate all possible choices for the sets $\mathcal{T}_{i}$. One possibility is to assign $\{u_{i},u_{i+1}\}\in\mathcal{I}(C^{c}_{7})$ to the vertices $v_{1}$ and $v_{2}$ and $\emptyset$ to the remaining vertices, which gives us $2w_{1}+2w_{2}\leq 1$. Similarly we have $2w_{i}+2w_{i+1}\leq 1$ for all $i=1,\dots,7$. Note that we use the convention $v_{i}=v_{j}$ iff $i=j~{}(\text{mod }7)$ here.
Now consider the sets $\{u_{i}\},\{u_{i},u_{i+1}\},\{u_{i+1}\},\emptyset$. We can assign $\{u_{i}\}$ to $v_{1}$, $\{u_{i},u_{i+1}\}$ to $v_{2}$ and $\{u_{i+1}\}$ to $v_{3}$ and $\emptyset$ to the remaining vertices, which gives us $w_{1}+2w_{2}+w_{3}\leq 1$ or more generally for all $i=1\dots 7$ we have $w_{i}+2w_{i+1}+w_{i+2}\leq 1$. Note that we can ignore these constraints, because adding $2w_{i}+2w_{i+1}\leq 1$ and $2w_{i+1}+2w_{i+2}\leq 1$ yields the same constraint.
Similarly for $4$ consequence vertices, $v_{i},v_{i+1},v_{i+2},v_{i+3}$ in $C_{7}$ we can assign $\{u_{1}\}$ to both $v_{i}$ and $v_{i+1}$ and $\{u_{2}\}$ to both $v_{i+2}$ and $v_{i+3}$, which gives us $w_{i}+w_{i+1}+w_{i+2}+w_{i+3}\leq 1$. Again, this inequality is redundant and can be obtained by $2w_{i}+2w_{i+1}\leq 1$ and $2w_{i+2}+2w_{i+3}\leq 1$. After carefully considering all possibilities for the sets $\mathcal{T}_{i}$, we only obtain the constraints $2w_{i}+2w_{i+1}\leq 1$, $i=1,\dots,7$. Then, we solve the linear program to obtain $\alpha^{*}(C_{7}|C_{7}^{c})=\frac{7}{4}$.
Appendix B
Haemers number, $\mathcal{H}(G)$, is an upper bound on the Shannon number of $G$, which considers the rank of particular matrices associated with the graph $G$. Let $G=C_{7}$ and $H=J(14,3)$. Since for any graph $G$, $\mathcal{H}(G)$ is an integer number, and $3.2578\leq\mathscr{C}(C_{7})$ [32] and $\mathscr{C}(C_{7})\leq\mathcal{H}(C_{7})$, we conclude that $\mathcal{H}(C_{7})\geq 4$. On the other hand in Section 3.1 we prove that $\mathcal{H}(J(14,3))=14$ and $\alpha^{*}(C_{7}|J(14,3))\leq\frac{1}{4}$. This implies that $\frac{\mathcal{H}(C_{7})}{\mathcal{H}(J(14,3))}\geq\frac{4}{14}>\frac{1}{4}\geq\alpha^{*}(C_{7}|J(14,3))$. So, the relative fractional independence number is not always an upper bound for the ratio of Haemers number of two given graphs.
Appendix C
We compute the value of $\alpha^{*}(C_{n}|C_{m})$. We would like to prove that when $n,m>1$, then
$\alpha^{*}(C_{n}|C_{m})=\begin{cases}\frac{n}{m-1}&\text{if $n$ is even and $m$ is odd,}\\ \frac{n}{m},&\text{if $m$ is even, }\\ \frac{n}{m},&\text{if $n$ is odd and $m$ is odd and $n\leq m$,}\\ \frac{n}{m-1},&\text{if $n$ is odd and $m$ is odd and $n>m$.}\par\end{cases}$
By Lemma 2,
$\alpha^{*}(C_{n}|C_{m})=\frac{n}{\alpha(C_{n}^{c}\boxtimes C_{m})}\geq\frac{\alpha^{*}(C_{n})}{\alpha^{*}(C_{m})}=\frac{n}{m}$
so $\alpha(C_{n}^{c}\boxtimes C_{m})\leq m$. Let $\{u_{1},\dots,u_{n}\}$ and $\{v_{1},\dots,v_{m}\}$ be the vertex sets of $C_{n}$ and $C_{m}$ respectively such that cyclically $u_{i}$ and $u_{i+1}$ and similarly $v_{j}$ and $v_{j+1}$ are connected. If $m\geq n$ and both are odd then we have a homomorphism $C_{m}\rightarrow C_{n}$ and hence by Lemma 5, $\alpha(C_{n}^{c}\boxtimes C_{m})\geq m$ and so $\alpha^{*}(C_{n}|C_{m})=\frac{n}{m}$. If $m$ is even then since $C_{m}$ is a perfect graph
$\alpha^{*}(C_{n}|C_{m})=\frac{\alpha^{*}(C_{n})}{\alpha(C_{m})}=\frac{n}{m}$
Similarly if $n$ is even and $m$ is odd then
$\alpha^{*}(C_{n}|C_{m})=\frac{\alpha(C_{n})}{\alpha(C_{m})}=\frac{n}{m-1}$
The only case remains when $n>m$, and they are both odd. In this case $(u_{i},v_{i})$ for $i=1,\dots,m-1$ is an independent set. So, we only need to show that we can not have $m$ independent vertices. Let $\mathcal{S}$ be a set of $m$ independent vertices of $C^{c}_{n}\boxtimes C_{m}$. There are two possibilities. Either, for each vertex $v_{k}$ of $C_{m}$, there is a vertex $(u_{i_{k}},v_{k})$ in $\mathcal{S}$, or there is a vertex, say without loss of generality, $v_{m-1}$ in $C_{m}$ that there is more than one vertex in $\mathcal{S}$ with second component equal to $v_{m-1}$. The first case can not happen since, in that case, $u_{i_{k}}$ is cyclically connected to $u_{i_{k+1}}$ but $n>m$, so it can not happen. In the second case, for any two connected vertices $\{a,b\}$ in $C_{m}$, there are at most $2$ vertices in $\mathcal{S}$ with the second component equal to $a$ or $b$. Since for any three vertices in $C_{n}$, at least two are not connected. Therefore, if there are three vertices in $\mathcal{S}$ with second component $a$ or $b$, one can find two vertices among them that are connected, a contradiction. Hence, if there is more than one vertex in $\mathcal{S}$ with a second component $v_{m-1}$, there must be exactly two such vertices, and there are no vertex in $\mathcal{S}$ with a second component equal to $v_{m}$ or $v_{m-2}$. Now for each pair $\{v_{2i-1},v_{2i}\}$ for $i=1,\dots,\frac{m-3}{2}$ one has at most $2$ vertex in $S$ with second component equal to $v_{2i-1}$ or $v_{2i}$. So we have at most $12(\frac{m-3}{2})+2=m-$ 1 vertices in $\mathcal{S}$. This shows that $\alpha(C_{n}^{c}\boxtimes C_{m})=m-1$ if $n>m$ and $n$ and $m$ are odd. Now, the calculation of $\alpha^{*}(C_{n}|C_{m})$ is complete.
Remark 7.
The best upper bound for the Shannon capacity of an odd cycle $C_{n}$ is Lovász number which is $\frac{n\cos{\pi/n}}{1+\cos{\pi/n}}=\dfrac{n}{2}-O\left(\dfrac{1}{n}\right)$[3]. A typical lower bound for the Shannon capacity of an odd cycle is via the independence number of its power graph. For example $\alpha(C^{2}_{m})=\left\lfloor{\dfrac{m^{2}-m}{4}}\right\rfloor$ and $\alpha(C^{d}_{m})\leq\dfrac{\alpha(C^{d-1}_{m})m}{2}$[33]. It is also known that when $m$ is of the form $k2^{d}+1$, then $\alpha(C^{d}_{m})={\left(\dfrac{m}{2}\right)}^{d}\cdot\dfrac{m-1}{m}$ [33]. Also, if $m=k2^{d}+1$ where $k$ is an odd positive number, then $\alpha(C^{d+1}_{m})\geq{\left(\dfrac{m}{2}\right)}^{d+1}-\dfrac{1}{2}\left({\left(\dfrac{m}{2}\right)}^{d}+m^{d-1}\right)$ (see [40]).
Using the lower bound of $\alpha(C^{2}_{m})=\left\lfloor{\dfrac{m^{2}-m}{4}}\right\rfloor$, we obtain
$\dfrac{\mathscr{C}(C_{n})}{\mathscr{C}(C_{m})}\leq\dfrac{\dfrac{n}{2}-O(\dfrac{1}{n})}{\sqrt{\left\lfloor{\dfrac{m^{2}-m}{4}}\right\rfloor}}.$
For large values of $n$, this bound is strictly worse than $n/m$. Similarly, one can verify that the bounds obtained using other lower bounds on $C_{m}$ are worse than $n/m$ for sufficiently large values of $n$. Therefore, the explicit upper bound
$\frac{\mathscr{C}(G)}{\mathscr{C}(H)}\leq\alpha^{*}(G|H)$ (37)
can be strictly better than the upper bound
$\frac{\mathscr{C}(G)}{\mathscr{C}(H)}\leq\frac{\mathscr{C}(G)}{(\alpha(H^{d}))^{1/d}}$ (38)
for those values of $d$ such that one can compute $\alpha(H^{d})$ explicitly.
Appendix D
In this appendix, we present an additional proof for Theorem 1 by building upon the proof of Hales in [8] for the fractional independence number. Additionally, we discuss the difference between our first proof and that of [8].
In [8], Hales proved that
$\alpha^{*}(G)=\max\sum_{v\in\mathcal{V}(G)}w_{v}=\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(W)}.$ (39)
Here, the maximum is taken over all weights $w_{v}\geq 0$ such that $\sum_{v\in\mathcal{S}}w_{v}\leq 1$ for every clique $\mathcal{S}\subset\mathcal{V}(G)$. The constraints in this linear program are defined based on the cliques of the graph $G$.
Defining the relative fractional independence number as $\alpha^{*}(G|H)=\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}$ is a natural extension of the fractional independence number. However, defining the linear function to maximize and the associated constraints when constructing a linear program model for computing $\alpha^{*}(G|H)$ is not immediately clear. Our first proof in Section 5.1 guides us to the correct form of the linear program. Nonetheless, if we have a guess for the form of the linear program, we can prove its correctness by extending the proof of [8].
Consider the linear program in part (i) of Theorem 1 where we maximize $\sum_{i=1}^{k}w_{i}$, subject to the constraints $\sum_{i=1}^{k}w_{i}|\mathcal{T}_{i}|\leq 1$ for any collection of sets $\mathcal{T}_{1},\cdots,\mathcal{T}_{k}\in\mathcal{I}(H)$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected in $H$ if there is no edge between $v_{i}$ and $v_{j}$ in $G$. Let $\eta(G|H)$ be the maximum of $\sum_{i=1}^{k}w_{i}$ in this linear program. We wish to show that $\eta(G|H)=\alpha^{*}(G|H)$ where
$\alpha^{*}(G|H)=\sup_{W}\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$
We first show that $\eta(G|H)\leq\alpha^{*}(G|H)$, and then show that $\eta(G|H)\geq\alpha^{*}(G|H)$.
Proof for $\eta(G|H)\geq\alpha^{*}(G|H)$: Take some arbitrary graph $W$. We need to show that
$\eta(G|H)\geq\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$ (40)
Let $\mathcal{S}$ be a maximum independent set for $G\boxtimes W$. For $v_{i}\in\mathcal{V}(G)$ define
$\mathcal{S}_{i}:=\{u\in\mathcal{V}(W):(v_{i},u)\in\mathcal{S}\}.$
Note that if $v_{i}$ and $v_{j}$ are adjacent, then $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ are disconnected.
Consider the assignment $w_{i}=\frac{|\mathcal{S}_{i}|}{\alpha(H\boxtimes W)}$. Note that
$\sum_{i}w_{i}=\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$
Thus, to show (40), it suffices to show that $w_{i}$’s satisfy the constraint of the linear program; that is for any subsets $\mathcal{T}_{1},\dots,\mathcal{T}_{n}$ of independent sets of $H$ such that $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected when there is no edge between $v_{i}$ and $v_{j}$ then
$\sum_{i=1}^{k}w_{i}|\mathcal{T}_{i}|\leq 1.$
Equivalently, we need to show that
$\sum_{i=1}^{k}|\mathcal{S}_{i}||\mathcal{T}_{i}|\leq\alpha(H\boxtimes W).$ (41)
First, note that the subsets $\mathcal{T}_{i}\times\mathcal{S}_{j}$ of $\mathcal{V}(H\boxtimes W)$ are pairwise disjoint. Indeed if $(x,y)$ belongs to $\mathcal{T}_{i}\times\mathcal{S}_{i}$ and $\mathcal{T}_{j}\times\mathcal{S}_{j}$ for $i\neq j$ then $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ are not disconnected and hence $v_{i}$ and $v_{j}$ are not adjacent and hence $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected (giving a contradicition). Next, observe that the union $\cup_{i=1}^{n}\mathcal{T}_{i}\times\mathcal{S}_{i}$ is an independent set of $H\boxtimes W$. To show this, take two vertices $(x_{1},y_{1})$ and $(x_{2},y_{2})$ in this set. If both of them belong to the same $\mathcal{T}_{i}\times\mathcal{S}_{j}$ then since $(v_{i},y_{1})$ and $(v_{i},y_{2})$ belong to the independent set $\mathcal{S}$ in $G\boxtimes W$ hence $y_{1}$ and $y_{2}$ are not adjacent in $W$. Therefore, the two vertices are not adjacent in $H\boxtimes W$.
Next, assume $(x_{1},y_{1})\in\mathcal{T}_{i}\times\mathcal{S}_{j}$ and $(x_{2},y_{2})\in\mathcal{T}_{j}\times\mathcal{S}_{j}$ and $i\neq j$. We are done if $y_{1}$ and $y_{2}$ are not adjacent. If $y_{1}=y_{2}$ then $v_{i}$ and $v_{j}$ are not adjacent and hence $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected. Therefore, $x_{1}$ and $x_{2}$ are not adjacent. Finally, if $y_{1}$ and $y_{2}$ are adjacent, then $v_{1}$ and $v_{2}$ are not adjacent and again $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ are disconnected. So we constructed an independent set of $H\boxtimes W$ of size $\sum_{v\in\mathcal{V}(G)}|S_{v}||T_{v}|$. So, (41) is proved, and the proof for this part is complete.
Proof for $\eta(G|H)\leq\alpha^{*}(G|H)$:
Let $w_{i}$ be any rational numbers satisfying the constraints of the linear program. We will construct a finite graph $W$ such that
$\sum_{i=1}^{k}w_{i}\leq\frac{\alpha(G\boxtimes W)}{\alpha(H\boxtimes W)}.$ (42)
This would establish that $\sum_{i=1}^{k}w_{i}\leq\alpha^{*}(G|H)$. Since the maximum of $\sum_{i=1}^{k}w_{i}$ over all the rational solutions is the same as this maximum over all real numbers, we will be done with this part.
It remains to give the construction of $W$. Let $N$ be the common denominator of $w_{1},\dots,w_{k}$. Let $W$ be a graph with $Nw_{1}+\dots+Nw_{k}$ vertices as follows: for each $v_{i}\in\mathcal{V}(G)$ take $\mathcal{D}_{i}$ be a set of vertices of $W$ consisting of $Nw_{i}$ vertices. There are no edges between the vertices in each $\mathcal{D}_{i}$. If $v_{i}$ and $v_{j}$ are not adjacent in $G$, then each vertex of $\mathcal{D}_{i}$ are connected by an edge to each vertex of $\mathcal{D}_{j}$; if $v_{i}$ and $v_{j}$ are adjacent in $G$ then there is no edge between any vertex of $\mathcal{D}_{i}$ and any vertex of $\mathcal{D}_{j}$. Thus, $W$ is essentially $G^{c}$ with the vertex $v_{i}$ being “copied” $Nw_{i}$ times.
Now, it is clear that the set
$\{(v_{i},y):v_{i}\in\mathcal{V}(G),y\in\mathcal{D}_{i}\}$
is an independent set in $G\boxtimes W$ of size $Nw_{1}+\dots+Nw_{k}$ and hence
$Nw_{1}+\dots+Nw_{k}\leq\alpha(G\boxtimes W).$ (43)
Take a maximal independent set $\mathcal{T}$ for $H\boxtimes W$ and for $y\in\mathcal{V}(W)$ let
$\mathcal{T}_{y}=\{u\in\mathcal{V}(H):(u,y)\in\mathcal{T}\}.$
Note that $\mathcal{T}_{y}$ is an independent set in $H$. Some observations are in order: Note that if $y_{1}$ and $y_{2}$ belong to $\mathcal{D}_{i}$ then $\mathcal{T}_{y_{1}}=\mathcal{T}_{y_{2}}$. Otherwise we can make $\mathcal{T}$ to a bigger independent set by adding $(u,y_{1})$ and $(u,y_{2})$ with $u\in\mathcal{T}_{y_{1}}\cup\mathcal{T}_{y_{2}}$ to $\mathcal{T}$. Since for $y_{1}$ and $y_{2}$ belonging to $\mathcal{D}_{i}$, we have $\mathcal{T}_{y_{1}}=\mathcal{T}_{y_{2}}$, we can use the notation $\mathcal{T}_{i}$ for $\mathcal{T}_{y}$ for any $y\in\mathcal{D}_{i}$. Therefore, assuming $\mathcal{V}(G)=\{1,2,\cdots,k\}$, we get the sets $\mathcal{T}_{1},\mathcal{T}_{2},\cdots,\mathcal{T}_{k}$.
If $v_{i}$ and $v_{j}$ are not connected in $G$, then vertices in the sets $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ are connected in $W$ and hence $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ must be disconnected in $H$ (otherwise $\mathcal{T}$ will not be an independent set for $H\boxtimes W$). Therefore, since the numbers $w_{1},\dots,w_{k}$ satisfy the constraints of our linear program, we get that $\sum_{i=1}^{k}w_{i}|\mathcal{T}_{i}|\leq 1$. Now since $|\mathcal{D}_{i}|=Nw_{i}$ we have
$\alpha(H\boxtimes W)=|\mathcal{T}|=\sum_{y\in\mathcal{V}(W)}|\mathcal{T}_{y}|=\sum_{i=1}^{k}Nw_{i}|\mathcal{T}_{i}|\leq N.$ (44)
From inequalities (43) and (44), the inequality (42) follows.
Remark 8.
Note that the construction of the graph $W$ depends on graph $H$ only through the weights $w_{i}$ of the linear program.
Appendix E
Proof of Theorem 9
Note that for a vertex-transitive graph $G$ of size $n$, $\alpha^{*}(G)=\frac{n}{\omega(G)}$, where $\omega(G)$ is the clique number of $G$. Thus, $\alpha^{*}(G)=\frac{n}{k+1}$ and $\alpha^{*}(H)=\frac{m}{k+1}$. It follows that $\alpha^{*}(G|H)\geq\frac{n}{m}$ and hence $\alpha(G^{c}\boxtimes H)\leq m$. If there is a homomorphism from $H$ to $G$ then by Lemma 5, we have $\alpha(G^{c}\boxtimes H)\geq m$ and hence we find that $\alpha^{*}(G|H)=\frac{n}{m}$. We show that such a homomorphism exists if there are integers $\ell,s\geq 0$ such that
$m=\ell n+s(k+1).$
Let $\{1,\dots,n\}$ and $\{1,\dots,m\}$ be the vertex sets for $G$ and $H$. If $\ell=0$ define $\mathsf{g}(i(k+1)+j)=j$ where $11\leq j\leq k+$ 1 and $i\geq 0$. If $\ell>0$, define $\mathsf{g}(in+j)=j$ for $1\leq j\leq n$ and $i=0,\dots,\ell-1$ and define $\mathsf{g}(\ell n+i(k+1)+j)=j$ for $11\leq j\leq k+$ 1 and $i\geq 0$. It can be verified that this is a homomorphism from $H$ to $G$.